### v1

# Monte Carlo Tree Diffusion with Multiple Experts for Protein Design

Abstract
The goal of protein design is to generate amino acid sequences that fold into functional structures with desired properties. Prior methods combining autoregressive language models with Monte Carlo Tree Search (MCTS) struggle with long-range dependencies and suffer from an impractically large search space. We propose MCTD-ME, Monte Carlo Tree Diffusion with Multiple Experts, which integrates masked diffusion models with tree search to enable multi-token planning and efficient exploration. Unlike autoregressive planners, MCTD-ME uses biophysical-fidelity-enhanced diffusion denoising as the rollout engine, jointly revising multiple positions and scaling to large sequence spaces. It further leverages experts of varying capacities to enrich exploration, guided by a pLDDT-based masking schedule that targets low-confidence regions while preserving reliable residues. We propose a novel multi-expert selection rule ($\text{P}\mathcal{H}\text{-UCT}\text{-ME}$) extends predictive-entropy UCT to expert ensembles. On the inverse folding task (CAMEO and PDB benchmarks), MCTD-ME outperforms single-expert and unguided baselines in both sequence recovery (AAR) and structural similarity (scTM), with gains increasing for longer proteins and benefiting from multi-expert guidance. More generally, the framework is model-agnostic and applicable beyond inverse folding, including de novo protein engineering and multi-objective molecular generation.
1 Introduction
Large generative models have demonstrated impressive capabilities across domains such as natural language processing [29] and molecular design [25]. Yet, conventional decoding strategies—greedy sampling, nucleus sampling, and beam search—remain fundamentally limited in tasks that require strategic planning or optimizing multiple objectives. In molecule or protein design, for example, a generative model must produce sequences that are not only syntactically valid but also meet biochemical, structural, and functional requirements [21]. Greedy decoding often gets stuck in suboptimal choices, and beam search—despite its popularity—offers little control over long-range dependencies or diverse trade-offs, especially in long-horizon domains like drug generation or code synthesis [3].
In contrast, planning algorithms such as Monte Carlo Tree Search (MCTS) offer adaptive search through a balance of exploration and exploitation. Since the introduction of UCT [11], MCTS has shown success in domains requiring sequential decision-making, most notably in AlphaGo [28]. Unlike static decoding, MCTS enables dynamic trajectory expansion, backtracking, and selective refinement, making it a promising mechanism to augment generative models with deliberative search. Recent frameworks such as Tree-of-Thoughts [33] and RethinkMCTS [13] illustrate this potential in language and code generation, where search paths can be evaluated and adjusted.
Large language models (LLMs such as GPT [18, 19, 22, 23, 24]) have recently been paired with MCTS to improve search in complex generation tasks. For instance, GPT-guided MCTS has enhanced search efficiency in symbolic regression [15], while value-guided MCTS decoding (PPO-MCTS [16]) improved the preferability of generated text over standard RLHF outputs. However, GPT-based MCTS approaches face notable pitfalls. Due to the autoregressive, token-by-token generation, even a single token error can derail an entire solution, and models often get stuck in suboptimal trajectories because their prior training biases lead to unbalanced exploration-exploitation. In other words, maintaining long-range coherence is challenging and it’s difficult to revise parts of a candidate without regenerating the whole sequence.
Beyond these GPT-specific pitfalls, integrating MCTS into large-scale generative frameworks—particularly for biological design—faces major challenges as follows: First, unlike classical planning environments, we lack a tractable forward model to evaluate partially generated sequences. While recent approaches like Diffuser [9] show how diffusion models can operate as planners without explicit forward simulators, standard diffusion decoding is still largely open-loop. This limits its ability to adaptively revise candidates during generation. Second, real-world biological design tasks are intrinsically multi-objective. Protein design, for instance, requires satisfying multiple criteria, from structural compatibility to folding efficiency and evolutionary plausibility. Third, GPT-based generators make it inefficient to apply MCTS, especially for long-sequence tasks, resulting in unnecessarily large tree searches in both breadth and depth and inevitably exploring unpromising states. Finally, relying on a single GPT for exploration limits biophysical-fidelity, diversity, and constrains the search space due to biases from the pretraining dataset.
In this paper, we propose Monte Carlo Tree Diffusion with Multiple Experts (MCTD-ME), a novel planning framework that integrates diffusion-based generation with an MCTS search guided by a diverse ensemble of domain experts models. Motivated by the challenges of protein sequence design—where search spaces are vastly larger than small-molecule generation—we leverage diffusion models for their ability to efficiently denoise multiple positions in parallel, and use expert models to narrow the search space by injecting biologically grounded priors. Each node in our tree represents a partially denoised sequence (a partial plan), and expansion corresponds to one step of noise reduction. This structure imbues the diffusion process with MCTS’s adaptive lookahead: evaluating, pruning, and refining paths to focus on promising candidates. In addition, we adopt an imitation-style rollout approach, where each expert proposes candidate completions based on the current partial sequence, thereby increasing diversity in the exploration space. These trajectories guide selection and backpropagation through the tree. Disagreement among experts encourages exploration; consensus promotes exploitation—enabling informed multi-objective planning at test time. We defer additional details on related work to Appendix 2.
Summary of contributions: To our knowledge, this is the first work to incorporate an ensemble of expert into a diffusion-based planning framework, enabling multi-experts Monte Carlo tree search for complex generative problems. By unifying diffusion models with a multi-critic evaluation module and a multi-expert proposal module within MCTS, we introduce a general test-time strategy to steer generation toward goals that single decoders or static samplers struggle with—especially in large, structured domains like protein sequence design, where generation spaces are vast and require both parallel decoding and multi-perspective exploration. Central to our design are (i) a pLDDT-guided masking policy that focuses edits on structurally uncertain regions to improve biophysical fidelity, and (ii) a multi-expert selection mechanism that arbitrates among experts using cached uncertainty statistics, yielding stable and efficient search.
Second, we propose several technical innovations to realize this integration:
  1. 1.
    A tree-structured diffusion process that expands and evaluates partial sequences through iterative biophysical-fidelity-enhanced denoising, achieved by masking low-pLDDT positions and guided by a multi-expert rollout mechanism, where each expert proposes candidate completions conditioned on the partial sequence, thereby enabling more diverse and meaningful exploration across a vast search space.
  2. 2.
    A novel $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$ selection algorithm, an entropy-aware selection rule, extending the classic UCT formula with exploration bonuses based on expert disagreement (e.g., score variance or entropy), inspired by entropy-guided UCB from ERP [21].
Third, we demonstrate empirical gains on long-horizon generation tasks—specifically inverse protein folding for just one case example. Using the CAMEO 2022 benchmark [10] and a PDB date-split benchmark [4], we show that our framework consistently improves sequence recovery (AAR) and structural similarity (scTM) over strong DPLM-2 baselines [31]. Our approach yields higher-quality, more diverse solutions, and these improvements grow with larger search budgets—highlighting the scalability and deliberative nature of our framework. We observe that both pLDDT-guided masking and multi-expert selection contribute complementary gains, with larger benefits on longer proteins where uncertainty is spatially concentrated. Overall, this work advances the state-of-the-art in test-time guided generative planning, opening up promising directions for applications in drug discovery, program synthesis, and beyond.
2 Related works
We review related work in protein diffusion models, MCTS in generative modeling and molecule design, discrete diffusion for planning and generation, and multi-objective and expert-guided Generation.
Protein and Drug Diffusion Models
Recent advances apply diffusion models to protein design across both structure and sequence domains. RFdiffusion [32] introduced a guided 3D diffusion model for protein backbones, achieving state-of-the-art results. On the sequence side, DPLM [30] and its extension DPLM-2 [31] demonstrated that discrete diffusion pretraining on large-scale protein data enables high-quality sequence generation and versatile conditioning (e.g., inverse folding, property steering). Most recently, DiffGui [8] proposed a target-aware, E(3)-equivariant diffusion framework for joint atom and bond generation, incorporating multi-property guidance (e.g., binding affinity, drug-likeness) to generate realistic, high-affinity 3D molecules. However, these diffusion approaches generally rely on one-shot or open-loop decoding, making it difficult to incorporate adaptive search or revision during generation—highlighting the need for planning methods such as MCTS.
Monte Carlo Tree Search in Generative Modeling
Monte Carlo Tree Search (MCTS) has improved generative decoding in tasks requiring complex reasoning or program synthesis. In code generation, AlphaCode [14] used a brute-force generate-and-test method by sampling large program sets and selecting those that passed tests. More structured methods followed: RethinkMCTS [13] performs MCTS over LLM reasoning steps, using execution feedback to refine intermediate thoughts. Tree-of-Thoughts [33] frames reasoning as a tree of self-evaluated partial solutions expanded via classical search (DFS/BFS), boosting performance on puzzles and planning. PG-TD [34] simulates lookahead in code generation by executing candidate programs during decoding. Across domains, MCTS-style planning consistently enhances generation quality and reliability. Yet, most of these approaches are tied to autoregressive GPT-style models, which remain token-by-token and struggle with long-range dependencies—suggesting that discrete diffusion could provide a more flexible backbone for planning.
Discrete Diffusion for Planning and Generation
While originally developed for continuous domains, diffusion models have been successfully adapted to discrete sequence generation and planning. D3PM [1] introduced absorbing-state (masking) noise for discrete diffusion, achieving competitive text generation. Follow-ups like [27] further closed the performance gap with autoregressive models. In planning, Diffuser [9] modeled offline RL as trajectory denoising, generating full state-action paths guided by value functions. To enable adaptive inference, Monte Carlo Tree Diffusion (MCTD) [2] reimagines diffusion as tree search, branching multiple denoising outcomes per step and scoring partial trajectories via heuristics. In protein design, ProtInvTree [17] similarly integrates reward-guided search with jumpy denoising for efficient inverse folding. Still, existing methods optimize toward a single diffusion generator, which limits their ability to effective search space exploration in biological and chemical design.
Multi-Experts Guided Generation
Multi-expert systems leverage several generative models or “experts” in tandem, allowing each to contribute from its area of strength. This idea has been explored in reinforcement learning and bandit settings, where algorithms learn from multiple expert policies or oracles to improve decision-making [20]. In generative modeling, mixture-of-experts strategies similarly combine specialized models to better cover complex data distributions. For example, recent diffusion-based image generators use multiple specialized denoising models across different noise levels [5] or even heterogeneous model architectures [12] to boost output quality. However, these multi-expert approaches remain largely static and one-shot during inference, without an adaptive search process. To our knowledge, no prior diffusion model integrates multiple experts into a planning loop for generative design. By incorporating a team of expert generators within an MCTS-driven diffusion framework, our work enables dynamic collaboration among experts to satisfy multiple design criteria and explore a broader solution space—an innovative combination that extends diffusion-based generation to more complex, multi-objective scenarios.
3 Preliminaries
We describe Diffusion-based generative model, MCTS, and protein design along with their mathematical notation in the subsequent sections and unify them under the Markov decision processes framework.
Markov decision processes.
We formulate generative planning as a finite-horizon MDP $\mathcal{M}=\langle\mathcal{S},\mathcal{A},H,P,R\rangle$ [26], where $\mathcal{S}$ is the set of states (e.g., partial or completed sequences), $\mathcal{A}$ is the action space (e.g., denoising steps or token insertions), and $P:\mathcal{S}\times\mathcal{A}\to\mathcal{S}$ defines stochastic transitions. The reward function $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ provides a scalar evaluation, often defined only on terminal states. The policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ defines the action distribution at each state, such as a denoising model in diffusion or a next-token model in autoregressive generation. An episode corresponds to constructing a sequence $y=(y_{1},...,y_{H})$, with a termination action marking the end of generation. We denote the action-value function $Q^{\pi}(s,a)$ as the expected cumulative reward of taking action $a$ in state $s$ under policy $\pi$, and the value function $V^{\pi}(s)$ as the expected return from $s$.
Diffusion-based generative model.
We briefly review the diffusion framework for sequence generation. Let $y_{0:H}$ denote a token sequence of length $H$. A forward noising process $q(y_{t}\mid y_{t-1})$, for $t=1,\ldots,T$, gradually corrupts the sequence, such that $y_{T}$ becomes maximally noisy (e.g., fully masked for discrete data or Gaussian noise for continuous). A reverse denoising model $p_{\phi}(y_{t-1}\mid y_{t})$ is trained to recover $y_{t-1}$ from $y_{t}$, enabling sampling via iterative denoising from $y_{T}\sim p(y_{T})$ down to $y_{0}\sim p_{\phi}(y_{0})$. Crucially, diffusion models naturally support guided generation. When an auxiliary score function $f(y)$ is available—e.g., a property predictor or constraint critic—generation can be biased toward desirable outputs. One classic approach is classifier guidance, where each reverse step is adjusted in the direction of $\nabla_{y_{t-1}}f(y_{t-1})$. Alternatively, reverse distribution can be reweighted as:
$p^{\text{guide}}_{\phi}(y_{t-1}\mid y_{t})\propto p_{\phi}(y_{t-1}\mid y_{t})\cdot\exp\left\{\beta\cdot f(y_{t-1})\right\},$
with $\beta$ controlling guidance strength. This formulation generalizes to multiple critics: given $K$ evaluators $C_{0},\dots,C_{K-1}$, each scoring complete sequences, guidance becomes even more natural. Instead of aggregating into a scalar reward, we perform imitation-style rollouts: partial candidates are completed, scored individually by each expert, and used to inform node evaluations during tree search. This preserves the diversity of expert preferences and supports structured planning, which we realize through our Monte Carlo Tree Diffusion (MCTD) framework.
MCTS.
Monte Carlo Tree Search (MCTS) is a planning algorithm that combines tree-based exploration with stochastic simulation to balance exploration and exploitation in large decision spaces. It proceeds in four stages: selection, expansion, simulation, and backpropagation. Starting from the root (current state), the selection phase recursively chooses child nodes using an Upper Confidence Bound (UCT) criterion [11]:
$\text{UCB}=Q\left({s,a}\right)+c_{p}\cdot\sqrt{\frac{\log\left({N\left({s}\right)}\right)}{N\left({s,a}\right)}},$ (1)
where $Q(s,a)$ is the estimated reward, $N(s)$ and $N(s,a)$ are visit counts, and $c_{p}$ controls exploration. When a leaf node is reached, expansion adds child nodes corresponding to unexplored actions. Then, simulation evaluates these new nodes by executing a rollout policy (e.g., sampling until termination or a predefined depth), producing a scalar reward. Finally, backpropagation propagates the reward up the visited path, updating $N(\cdot)$ and $Q(\cdot)$ statistics. Over repeated iterations, MCTS refines its value estimates and focuses the search on high-reward regions of the state space.
4 The MCTD-ME Algorithm
Figure 1: Overview of MCTD in diffusion-based protein sequence generation. Nodes are partially denoised sequences expanded via masked diffusion and guided by multi-expert rollouts.
We propose Monte Carlo Tree Diffusion with Multi-Experts (MCTD-ME), which enhances protein generation by combining masked diffusion with an MCTS planner. MCTD-ME structures diffusion as a tree, refining partially masked sequences to preserve scaffold integrity while systematically exploring uncertain regions. Search quality and diversity are further improved by incorporating multiple diffusion experts for diverse refinements and by using pLDDT-guided masking to target low-confidence residues. These design choices couple the planning strength of tree search with expert-aware diversification, yielding a principled framework for balancing exploration and exploitation in protein modeling. We present the algorithm in Algorithm 1, with each rollout consisting of four key steps.
Selection.
To navigate the space of protein sequences under the reverse masked diffusion framework, we define a tree where each node corresponds to a complete sequence (with no masked tokens) at a particular reverse step, and edges represent one reverse transition obtained by applying masked discrete diffusion to a low-confidence subset of tokens. At each iteration, we employ a novel selection algorithm, $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$—an entropy-augmented variant of UCT—designed to identify the most promising path, where ME denotes Multiple Experts. Specifically, from a parent node $s_{t}$ we choose the transition (i.e., reverse move to $s_{t-1}$ via a particular mask set) that maximizes the following objective:
$\text{P}\mathcal{H}\text{-UCT}\text{-ME}(s_{t})\;=\;\mathop{\rm arg\,max}_{a\in\mathcal{A}}\;\text{P}\mathcal{H}\text{-UCB}\text{-ME}\!\left((s_{t},a)\right),$ (2)
with
$\text{P}\mathcal{H}\text{-UCB}\text{-ME}\!\left((s_{t},a)\right)$ $=Q(s_{t},a)+c_{p}\,\frac{\sqrt{\log N(s_{t})}}{1+N(s_{t},a)}\cdot\left({\underbrace{w_{\mathrm{ent}}\,\mathcal{U}_{\text{ent}}(s_{t},a)\;+\;w_{\mathrm{div}}\,\mathcal{U}_{\text{div}}(s_{t},a)}_{\text{$\text{P}\mathcal{H}$\text{-ME} bonus (cached at expansion)}}}\right).$ (3)
Here, $Q(s_{t},a)$ is the current value estimate, $N(\cdot)$ are visit counts, $c_{p}>0$ controls exploration, and $w_{\mathrm{ent}},w_{\mathrm{div}}\!\geq\!0$ weight the predictive-uncertainty and diversity bonuses, respectively. In our diffusion setting, the action prior is given by the reverse diffusion model $p_{\phi}(s_{t-1}\!\mid\!s_{t})$ over masked edits, but—consistent with our implementation—we do not recompute its entropy during selection; instead we cache uncertainty statistics for each expanded child and reuse them during subsequent selections.
Multi-expert PH bonus. Let $\mathcal{M}$ be the mask set chosen at $s_{t}$ (e.g., via pLDDT), and let $\{p_{\phi}^{(e)}\}_{e=1}^{E}$ denote the $E$ expert conditional distributions (softmax of logits) used during expansion. For the candidate child $s_{t-1}$ obtained by action $a$, we define the ensemble-based uncertainty as
$\mathcal{U}_{\text{ent}}(s_{t},a)=\mathcal{H}\!\left(\frac{1}{E}\sum_{e=1}^{E}p_{\phi}^{(e)}(\,\cdot\,\mid s_{t-1},\mathcal{M})\right)-\frac{1}{E}\sum_{e=1}^{E}\,\mathcal{H}\!\left(p_{\phi}^{(e)}(\,\cdot\,\mid s_{t-1},\mathcal{M})\right).$ (4)
Here $\mathcal{H}[p]=-\sum_{a}p(a)\log p(a)$ denotes Shannon entropy. The first term is the entropy of the ensemble-averaged distribution $\bar{p}=\tfrac{1}{E}\sum_{e}p^{(e)}$, while the second is the average entropy of the individual experts. Their difference equals the mutual information (BALD score [7]), or Jensen–Shannon–style divergence between experts. Intuitively, it isolates epistemic uncertainty: if experts disagree, $\mathcal{H}[\bar{p}]$ is large while the average $\mathcal{H}[p^{(e)}]$ remains small, giving a positive signal. If all experts are merely noisy but agree, the two terms cancel.
We also include a diversity measure, defined as the normalized Hamming distance between parent and child sequences:
$\mathcal{U}_{\text{div}}(s_{t},a)$ $=\frac{1}{L}\sum_{i=1}^{L}\,\mathbf{1}\{y^{\text{child}}_{i}\neq y^{\text{parent}}_{i}\},$ (5)
For our inverse folding case studies, this simply counts the fraction of amino acids that differ. Both $\mathcal{U}_{\text{ent}}$ and $\mathcal{U}_{\text{div}}$ are computed once at expansion and stored for reuse during selection.
An action $a$ is prioritized if it (i) has high estimated value $Q(s_{t},a)$, (ii) corresponds to an under-explored branch (UCB exploration), (iii) induces strong ensemble disagreement ($\mathcal{U}_{\text{ent}}$ large), indicating information gain, and (iv) introduces sufficient novelty relative to the parent ($\mathcal{U}_{\text{div}}$ large). This encourages the search to balance reward exploitation with uncertainty- and diversity-driven exploration, while the reverse diffusion prior $p_{\phi}(s_{t-1}\!\mid\!s_{t})$ maintains consistency with the denoising dynamics.
Figure 2: Masked discrete diffusion step: low-confidence sites (mask) are resampled while confident tokens remain fixed, yielding a progressively “unmasked” sequence.
Expansion.
When $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$ selects a leaf node $s_{t}$, representing a complete sequence $y_{t}$, we expand it by generating candidate children via masked diffusion. Our expansion departs from standard MCTS in two important ways: (1) structure-aware masking via pLDDT scores to enhance biophysical fidelity, and (2) multi-expert imitation-guided rollouts to improve diversity and quality.
Progressive pLDDT Masking for Targeted Denoising. Rather than resampling all positions uniformly, we exploit predicted structural confidence (pLDDT) to target only low-confidence residues. Specifically, at reverse step $t$ we define a mask $\mathcal{M}_{t}$ by thresholding the pLDDT profile of $y_{t}$, masking unstable regions while preserving high-confidence subsequences. Crucially, the masking threshold decreases over diffusion steps, progressively reducing the fraction of masked residues as denoising proceeds. This “progressive masking” strategy focuses computation on the most uncertain regions early on, while stabilizing promising motifs in later stages:
$y_{t-1}\sim p_{\phi}(y_{t-1}\mid\text{Mask}(y_{t};\mathcal{M}_{t})).$ (6)
This masked input encourages structural refinement and prevents overwriting already stable subsequences, reducing unnecessary sampling noise and improving convergence.
Multi-Expert Guided Expansion via Rollouts. To enrich proposal diversity and better imitate the training distribution, we leverage multiple pretrained experts during expansion. Given a masked sequence $\text{Mask}(y_{t};\mathcal{M}_{t})$, each expert $e\in\{1,\dots,E\}$ performs $k$ rollouts by sampling its conditional distribution $p_{\phi}^{(e)}(\cdot\mid y_{t},\mathcal{M}_{t})$. This yields a pooled candidate set
$\mathcal{C}_{t}\;=\;\bigcup_{e=1}^{E}\ \bigcup_{r=1}^{k}\ y_{t-1}^{(e,r)},$ (7)
where $y_{t-1}^{(e,r)}$ denotes a complete sequence obtained by filling the masked sites and splicing them back into $y_{t}$. All candidates in $\mathcal{C}_{t}$ are de-duplicated and scored once with the composite reward (see Evaluation). We then retain the top-$K$ children, each annotated with two exploration bonuses computed at expansion: an ensemble surprisal term $\mathcal{U}_{\text{ent}}$ (capturing expert disagreement on the masked sites) and a novelty term $\mathcal{U}_{\text{div}}$ (the normalized Hamming distance to the parent). These cached quantities are reused during $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$  selection, ensuring that search is guided by both expert-informed proposal diversity and explicit uncertainty signals.
Candidate Proposal Strategy. To retain diversity while prioritizing promising edits, we rank all expanded children and select the top-$K$ candidates:
$\left\{y_{i=1}^{K}\right\}=\operatorname{Top\text{-}K}\big(\{y_{t-1}^{(j)}\},K\big).$ (8)
Here, $j$ indexes every candidate child pooled across experts and rollouts at step $t$; $\operatorname{Top\text{-}K}$ ranks them by the composite score and returns the best $K$. Each chosen child is inserted into the tree as a new node, annotated with its cached $\text{P}\mathcal{H}\text{-ME}$ bonuses ($\mathcal{U}_{\text{ent}},\mathcal{U}_{\text{div}}$) and expert rollout scores. This beam-style expansion ensures that multiple high-quality hypotheses are pursued in parallel, enabling MCTD-ME to balance exploitation of strong candidates with exploration of diverse structural alternatives.
Evaluation.
Unlike expansion, where multiple experts serve as proposal generators, the evaluation stage relies on a set of critics to score complete sequences. Because every node in the tree corresponds to a fully denoised sequence $y_{0}$, thus evaluation can be performed immediately without further decoding. Each sequence is scored by a panel of critics $\{C_{1},\dots,C_{J}\}$, producing values
$C_{j}(y_{0})\;:=\;R_{j}(y_{0}),\quad j=1,\dots,J,$ (9)
where the $C_{j}$ represent evaluation functions such as sequence recovery (AAR), structural similarity (scTM), or biophysical quality metrics, $J$ here denotes the number of critics. In practice, we aggregate these scores into a single scalar via a fixed weighted sum
$R(y_{0})\;=\;\sum_{j=1}^{J}w_{j}\,C_{j}(y_{0}),\qquad w_{j}\geq 0,\ \sum_{j=1}^{J}w_{j}=1,$ (10)
This design matches the diffusion context, where evaluation is only meaningful at complete denoised sequences.
Backpropagation.
The expert rewards obtained at $s_{t-1}$ are propagated back to update values along the path to the root. For each state–action pair $(s_{i},a_{i})$ on the path, we update:
$Q(s_{i},a_{i})\;\leftarrow\;\max\{Q(s_{i},a_{i}),\;r_{H}\},$ (11)
where $r_{H}$ is an aggregated expert score (e.g., the max across $\{R_{e}\}$). This max-based backup, consistent with our implementation, favors exploiting high-reward branches while maintaining the ability to explore alternative completions. Because each node is a complete sequence, long-range credit assignment is naturally aligned with the denoising trajectory.
Variant AAR Normalized Reward scTM
CAMEO Benchmark
Baseline Final $\Delta$ Baseline Final $\Delta$ Baseline Final $\Delta$
MCTD-ME-0 (Random) 0.438 0.381 0.057-0.057 0.425 0.429 +0.004+0.004 0.362 0.458 +0.096+0.096
Single-Expert (150M, Exp. 1) 0.427 0.427 +0.000+0.000 0.290 0.290 +0.000+0.000 0.371 0.371 +0.000+0.000
Single-Expert (650M, Exp. 0) 0.442 0.458 +0.016+0.016 0.431 0.447 +0.016+0.016 0.372 0.387 +0.014+0.014
Single-Expert (3B, Exp. 2) 0.443 0.429 0.014-0.014 0.432 0.422 0.011-0.011 0.373 0.369 0.005-0.005
MCTD-ME (Multi-Expert) 0.439 0.451 +0.012 0.427 0.465 +0.038 0.365 0.449 +0.084
PDB Benchmark
MCTD-ME-0 (Random) 0.430 0.400 0.030-0.030 0.110 0.100 0.010-0.010 0.360 0.380 +0.020+0.020
MCTD-ME-1 (650M) 0.490 0.520 +0.030+0.030 0.330 0.360 +0.030+0.030 0.390 0.420 +0.030+0.030
MCTD-ME-1 (150M) 0.460 0.460 +0.000+0.000 0.290 0.290 +0.000+0.000 0.370 0.370 +0.000+0.000
MCTD-ME-1 (3B) 0.500 0.420 0.080-0.080 0.340 0.280 0.060-0.060 0.400 0.370 0.030-0.030
MCTD-ME (Multi-Expert) 0.500 0.560 +0.060 0.350 0.410 +0.060 0.880 0.910 +0.030
Table 1: Main results on CAMEO and PDB. Baseline and final values for AAR, normalized reward, and scTM are shown, with $\Delta$ indicating improvements. Multi-expert MCTD consistently outperforms single-expert and unguided baselines, especially on PDB targets.
5 Experiments
5.1 Protein Inverse Folding
We evaluate Monte Carlo Tree Diffusion with Multiple Experts (MCTD-ME) on the task of protein inverse folding, i.e. structure-conditioned sequence design. The main purpose of inverse folding is to find the amino acid sequence such that its folding matches a given backbone structure. In this work, we focus on this domain where ground truth sequences are known, leaving other applications (e.g. de novo structure generation) for future work. The evaluation uses the CAMEO 2022 dataset [10] and PDB data split ([4]), a standard benchmark of recently solved protein structures used in prior work. For each backbone structure in CAMEO, the goal is to generate an amino acid sequence that folds to that structure. We adopt the same metrics as DPLM-2 [31]: (1) Amino Acid Recovery (AAR) – the percentage of residues in the designed sequence that exactly match the native sequence (sequence recovery), and (2) self-consistency TM-score (scTM) – the TM-score measuring structural consistency between the target backbone and the structure predicted from the designed sequence. Higher AAR indicates better sequence accuracy for the given fold, while scTM reflects how well the designed sequence is likely to fold into the target structure. Following DPLM-2’s protocol, we report both metrics for the CAMEO test set. (For reference, co-generation baselines like MultiFlow [4] and ESM3 [6] attain  32–47% AAR on CAMEO, while the state-of-the-art diffusion model DPLM-2 (3B) achieves  52% AAR and 0.89 scTM)
5.1.1 Baselines and Variants
We conduct an ablation study to quantify the impact of multi-expert guidance. In particular, we compare three variants of our approach:
MCTD-ME-0 (no experts): A control variant that performs the diffusion tree search with no expert guidance. The forward diffusion model’s proposals are accepted at random, yielding essentially a random-fill of the sequence (no reward-driven pruning).
MCTD-ME-1 (single expert): Our MCTS-based diffusion planner using a single expert model as the critic. Here we use the DPLM-2 150M, 650M, 3B models seperately as the sole experts for scoring candidates (chosen for its strong performance and manageable runtime). This represents a single-criterion search.
MCTD-ME-E (ensemble of experts): The full multi-expert MCTD using an ensemble of three DPLM-2 models of varying scales – 150M, 650M, and 3B parameters. These experts provide a diverse set of evaluations, which are combined into a composite reward signal.
All variants use the same MCTS sampling pipeline described in Section 4. At each diffusion step (node expansion), the planner simulates rollouts to incrementally denoise the sequence. In our implementation, each expert performs 2 rollouts, and with 3 experts this yields around 6 candidate expansions per node. To limit branching, we retain the Top-2 candidates according to the multi-critic reward at each step. The MCTS selection policy uses a UCT-style formula adapted for multi-expert value estimates with an exploration constant $c_{p}=1.414$. We also incorporate an entropy-aware bonus as described (Section 4.3), so that nodes with high disagreement among experts are explored more. All expert models are pre-trained (no further fine-tuning); MCTD operates as a test-time algorithm that plugs these experts into the search.
Reward Formulation: The multi-objective reward $R$ used to guide MCTD-E is a weighted sum of the expert criteria. Specifically, we define $R=0.6\text{AAR}+0.35\text{scTM}+0.05B$, where B is a small bonus from simple biophysical properties (average hydrophobicity, amino acid composition frequency, and net charge). The 60/35/5 weighting was chosen to prioritize sequence recovery while still strongly rewarding structural fidelity, plus a minor regularization to favor realistic amino acid distributions. This composite reward design is in line with prior multi-critic strategies. In our setting, AAR and scTM are computed after a full sequence is generated (i.e. at a leaf node of the search); partial sequences during tree expansion receive interim estimates (for example, we estimate scTM for a partial sequence by filling missing residues with the most likely amino acids before structure prediction, and likewise compute a partial AAR by comparing the filled-in sequence to the native). These approximations guide the search towards promising candidates even before the sequence is complete.
5.1.2 Results
Table 1 summarizes the inverse folding results on the CAMEO dataset. Overall, MCTD-ME-E (multi-expert) achieves the best performance, outperforming both the single-expert and no-expert variants by a significant margin. MCTD-ME-E attains an average AAR of 44.5%, higher than the 43.2% achieved by the strongest single-expert run (650M, expert 0). The normalized reward of MCTD-ME-E designs averages 0.303, also exceeding the 0.295 of the 650M expert alone. By contrast, the MCTD-ME-1 (150M expert, expert 1) variant essentially matches its baseline (42.7% AAR, 0.290 reward), while the MCTD-ME-1 (3B expert, expert 2) variant actually degrades performance (37.0% AAR, 0.254 reward). This indicates that our MCTS planner by itself introduces little overhead or degradation when only one expert is used – essentially reproducing the single model’s capabilities – but the ensemble of experts (MCTD-ME-E) yields a further boost: +1.3 percentage points in sequence recovery and a notable gain in normalized reward relative to the best single expert. This suggests that the ensemble is successfully steering the generation to find sequences that satisfy multiple criteria better than any single model alone. As expected, MCTD-ME-0 (random fill) performs near chance-level – we observed 38.7% AAR (roughly the random identity baseline for 20 amino acids) and very low reward (0.272), confirming that unguided diffusion yields essentially arbitrary sequences with poor folding to the target. These serve as a sanity check that improvements of MCTD-ME-1 and MCTD-ME-E are indeed due to learned guidance rather than artifact.
Importantly, MCTD-ME-E not only improves average performance but also reduces failure cases. We found that MCTD-ME-E produced valid, foldable sequences for nearly all test structures, whereas single-expert runs occasionally returned sequences with poor recovery or structural inconsistencies. The multi-expert approach mitigates this by combining perspectives: the smaller 150M expert, for instance, tends to propose more diverse amino acid choices (higher entropy), while the larger 3B expert strongly evaluates overall fitness. The MCTS search benefits from this diversity–exploitation mix, as reflected in the higher overall reward of MCTD-ME-E designs. We also observed that the advantage of multiple experts grows with problem complexity. On longer proteins (length >300), MCTD-ME-E’s sequence recovery was several percentage points higher than MCTD-ME-1 on average, whereas on short proteins (<100 residues) both methods performed similarly. This trend aligns with the intuition that larger search spaces and more complex structures gain more from an ensemble of critics guiding the search. In sum, multi-expert MCTD consistently outperforms both single-expert planning and one-shot diffusion decoding on this task, achieving state-of-the-art results on the CAMEO inverse folding benchmark (to our knowledge). See Fig 3 for length-binned improvements in AAR, reward, and scTM.
Figure 3: Comparison of three evaluation metrics (AAR, normalized reward, and scTM) across different model variants. Higher values indicate better sequence recovery and structural consistency, showing the advantage of multi-expert MCTD.
6 CONCLUSION
In conclusion, we introduced MCTD-ME, which combines diffusion-based generative planning with MCTS and multi-expert guidance. The method is novel in reframing diffusion decoding as tree search for parallel multi-token exploration and in integrating multiple expert evaluators with a pLDDT-guided masking schedule to improve biophysical fidelity. Our selection policy ($\text{P}\mathcal{H}\text{-UCT}\text{-ME}$) efficiently arbitrates among experts using cached uncertainty statistics, enabling stable and low-overhead rollouts. Empirically, MCTD-ME outperforms state-of-the-art baselines on inverse folding, scales effectively with additional computation, and provides a general template for diffusion-plus-search workflows in structured generative tasks.
Appendix A Appendix
A.1 Algorithm details
Algorithm 1 Monte Carlo Tree Diffusion with Multiple Experts (MCTD-ME)
1:
[…]
4:$K$: number of children per expansion
5:$E$: number of expert models
6:$R$: rollouts per expert
7:$T$: total simulations
[…]
11:for $i=1$ to $T$ do
12:  $node\leftarrow root$
13:  $\triangleright$ /* Selection via $\text{P}\mathcal{H}\text{-UCT}\text{-ME}*$/
14:  while $node.children\neq\emptyset$ do
15:   $node\leftarrow\text{P}\mathcal{H}\text{-UCT}\text{-ME}(node.children)$ // uses $Q$, UCB, and cached $(\mathcal{U}_{\text{ent}},\mathcal{U}_{\text{div}})$
16:  $\triangleright$ /* Expansion via pLDDT Masking and Multi-Expert Rollouts */
17:  $\mathcal{M}\leftarrow\textsc{GetMaskSet}(node.sequence)$ // Based on pLDDT
18:  $\mathcal{Y}\leftarrow\emptyset$ // candidate children
19:  for $e\in\mathcal{E}$ do // each generator expert
20:   for $r=1$ to $R$ do // rollouts
21:     $y_{0}\leftarrow\textsc{MaskedDiffusion}(node.sequence,\mathcal{M},e)$
22:     if $y_{0}\notin cache$ then
23:      $score\leftarrow\textsc{EvalComposite}(y_{0};\mathcal{C})$ // weighted sum of critics
24:      $cache[y_{0}]\leftarrow score$
25:     $\mathcal{Y}\leftarrow\mathcal{Y}\cup\{y_{0}\}$
26:  $\mathcal{Y}_{\text{top}}\leftarrow\textsc{TopK}(\mathcal{Y},K;\,cache)$ // sort by composite reward
27:  for $y^{\prime}\in\mathcal{Y}_{\text{top}}$ do
28:   $(\mathcal{U}_{\text{ent}}(y^{\prime}),\mathcal{U}_{\text{div}}(y^{\prime}))\leftarrow\textsc{ComputeBonuses}(y^{\prime},\,node.sequence,\,\mathcal{M})$ // ensemble surprisal & novelty
29:   $\textsc{AddChild}(node,y^{\prime},\,\mathcal{U}_{\text{ent}}(y^{\prime}),\,\mathcal{U}_{\text{div}}(y^{\prime}))$ // create new node with full seq
30:  $\triangleright$ /* Evaluation of New Children (full-sequence) */
31:  for $y^{\prime}\in\mathcal{Y}_{\text{top}}$ do
32:   $v\leftarrow cache[y^{\prime}]$ // may hit cache; same composite reward as above
33:   $\triangleright$ /* Backpropagation */
34:   $\textsc{Backpropagate}(y^{\prime},\,v)$ // supports max or sum backup
35:return Top-$k$ sequences by $cache[\cdot]$
A.2 Dataset Details
We evaluate MCTD-ME on two protein sequence benchmarks at the chain level. (1) CAMEO2022 comprises 183 targets released in 2022; we use the reference backbones and evaluate inverse folding on the corresponding chains. (2) PDB_date is a date-split benchmark of 7,855 proteins constructed from the Protein Data Bank (PDB); we hold out targets by deposition date and evaluate inverse folding on their chains. In our experiments we use a date-split subset of 449 chains from PDB_date for compute efficiency; unless otherwise noted, length statistics below are for the full benchmark. For both sets, sequence lengths are computed after simple canonicalization (upper-casing and removal of non-canonical/unknown residues).
Dataset # Proteins Min Mean Max
CAMEO2022 183 15 247.5 704
PDB_date (full) 7,855 2 242.2 511
Table 2: Length statistics (in residues) of amino-acid sequences in each benchmark. Counts refer to the number of evaluated chains. Note: Experiments use a date-split subset of 449 PDB_date chains; its length distribution is similar to the full set.
A.3 Case Study on CAMEO data
Fig. 4 overlays three structures to illustrate how MCTD-ME steers sequences toward the target fold. The red ribbon is the CAMEO reference (ground truth), the blue ribbon is the structure predicted from the baseline sequence, and the multicolor ribbon is our MCTD-ME design. For the multicolor model, residues are shaded on a red $\rightarrow$ blue gradient proportional to per-residue deviation from the reference (red = near, blue = far), making local improvements visually apparent along the chain. Relative to the blue baseline, the MCTD-ME design shows widespread “reddening,” especially across helices and connecting loops, indicating closer alignment to the target geometry while preserving already well-formed secondary structure. This behavior matches the algorithm’s intent: pLDDT-aware masking focuses edits on uncertain segments, while multi-expert rollouts—selected via $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$—propose diverse fixes that are promoted during search, yielding higher structural fidelity (e.g., scTM) without over-editing stable regions.
Figure 4: CAMEO 7dz2_C overlay. Red: target reference. Blue: baseline (starting point). Multicolor: MCTD design, colored per residue from red (near reference) to blue (far), after structural alignment. The shift from blue toward red across many regions visualizes how MCTD moves the baseline toward the reference via pLDDT-aware masking and multi-expert guided rollouts.
A.4 Setup and Hyperparameters
For equitable assessment, we compare all variants under matched search and decoding budgets. Unless otherwise specified, every method uses the same DPLM-2 decoding configuration (deterministic unmasking with argmax sampling) and identical reward evaluation. In each table/figure, competing methods share the same rollout budget and selection rule; only the availability of diffusion experts differs across modes (random_no_expert, single_expert, multi_expert). We conducted only light tuning around a small grid and then fixed a single setting for all reported comparisons.
Component Symbol Setting(s) used
Rollouts per expert $k_{\text{roll}}$ 3 (fixed for reported runs)
Children kept per expansion $K$ 3 (beam size per expansion)
Forward steps (tree depth) max_depth 5 (default; held fixed unless noted)
Exploration constant ($\text{P}\mathcal{H}\text{-UCT}\text{-ME}$) $c_{p}$ 1.414 (UCB1 constant)
Candidates per expansion (random mode) $N_{\text{cand}}$ 6 (used only in random_no_expert)
Backup rule max (value backup along the path)
Number of experts (multi_expert) $E$ 3 (unless otherwise noted)
Modes compared random_no_expert,  single_expert,  multi_expert
Single-expert ID in plots single_expert_0 (for direct comparisons)
Diffusion reverse steps max_iter 150
Diffusion temperature $T$ 1.0
Decoding strategy (all modes) deterministic unmasking,  argmax sampling
Table 3: Search setup and hyperparameters. We keep decoding and evaluation identical across methods. Hyperparameters are fixed for the main comparisons; only the presence/number of experts differs by mode.
A.5 Runtime and Efficiency
The enhanced performance comes at the cost of additional computation. MCTD-ME must evaluate multiple experts across many tree expansions, whereas a single diffusion model decoding is much faster. In our current implementation, generating one sequence for a CAMEO structure with MCTD-ME-E takes on the order of a few minutes (wall-clock) on a single GPU – roughly an order of magnitude slower than a single DPLM-2 pass. We emphasize that our focus here is on proof-of-concept effectiveness; optimization of the search procedure is left for future work. Techniques such as parallelizing rollouts, caching expert evaluations, or pruning low-reward branches more aggressively could greatly speed up MCTD-ME without sacrificing quality.
Experiments were run on a SLURM-managed cluster with one NVIDIA A100 GPU per job, four CPU cores and 32 GB RAM. All results were obtained with PyTorch mixed precision and identical inference code across modes; wall-time numbers in Table 4 are measured end-to-end per target (including expert rollouts and critic scoring).
Mode Min (s) Mean (s) Max (s)
Random (no expert) 89.6 645.3 5235.6
Single-expert (650M) 154.9 345.7 1024.8
Single-expert (150M) 257.6 502.9 1791.6
Single-expert (3B) 148.6 451.7 1052.0
[…]
Table 4: Per-target wall-clock time (seconds) aggregated over CAMEO2022 and PDB_date runs. “Single-expert (650M/150M/3B)” correspond to our three DPLM-2 capacities (IDs 0/1/2). Times include all steps (masking, diffusion rollouts, scoring, and tree updates).
A.6 Generality
Although our evaluation focused on the inverse folding problem, the MCTD-ME framework is inherently general and model-agnostic. Any diffusion-based generator can be plugged into the tree search, and any collection of expert predictors can shape the reward. This flexibility opens up many possibilities: for instance, MCTD-ME could be applied to de novo protein design by incorporating structural-validity scores and evolutionary metrics as critics, or extended to other domains such as molecule generation using pharmacophoric and ADMET property predictors. Our results on protein inverse folding demonstrate the promise of combining diffusion models with MCTS guided by multiple critics for multi-objective generative design. We believe this diffusion+MCTS approach offers a powerful new paradigm for generative modeling, and extending MCTD-ME to more tasks with richer reward functions is an exciting direction for future work.

### v2

@@
 Prior methods combining autoregressive language models with Monte Carlo Tree Search (MCTS) struggle with long-range dependencies and suffer from an impractically large search space.
-We propose MCTD-ME, Monte Carlo Tree Diffusion with Multiple Experts, which integrates masked diffusion models with tree search to enable multi-token planning and efficient exploration.
+We propose MCTD-ME, Monte Carlo Tree Diffusion with Multiple Experts, which integrates masked diffusion models with tree search to enable multi-token planning and efficient exploration under the guidance of multiple experts.
 Unlike autoregressive planners, MCTD-ME uses biophysical-fidelity-enhanced diffusion denoising as the rollout engine, jointly revising multiple positions and scaling to large sequence spaces.
 It further leverages experts of varying capacities to enrich exploration, guided by a pLDDT-based masking schedule that targets low-confidence regions while preserving reliable residues.
-We propose a novel multi-expert selection rule ($\text{P}\mathcal{H}\text{-UCT}\text{-ME}$) extends predictive-entropy UCT to expert ensembles.
-On the inverse folding task (CAMEO and PDB benchmarks), MCTD-ME outperforms single-expert and unguided baselines in both sequence recovery (AAR) and structural similarity (scTM), with gains increasing for longer proteins and benefiting from multi-expert guidance.
-More generally, the framework is model-agnostic and applicable beyond inverse folding, including de novo protein engineering and multi-objective molecular generation.
+We propose a novel multi-expert selection rule ($\text{P}\mathcal{H}\text{-UCT}\text{-ME}$) extends Shannon-entropy-based UCT to expert ensembles with mutual information.
+MCTD-ME achieves superior performance on the CAMEO and PDB benchmarks, excelling in protein design tasks such as inverse folding, folding, and conditional design challenges like motif scaffolding on lead optimization tasks.
+Our framework is model-agnostic, plug-and-play, and extensible to de novo protein engineering and beyond.
+Machine Learning, ICML
+[…]
+$\Statex\ALG@thistlm\triangleright #1 \algnewcommand\IndentLineComment$$\Statex\ALG@tlm\triangleright #1$
 1 Introduction
@@
 Recent frameworks such as Tree-of-Thoughts [40] and RethinkMCTS [15] illustrate this potential in language and code generation, where search paths can be evaluated and adjusted.
-Large language models (LLMs such as GPT [18, 19, 22, 23, 24]) have recently been paired with MCTS to improve search in complex generation tasks.
+Large language models (LLMs such as GPT [25, 26, 27]) have recently been paired with MCTS to improve search in complex generation tasks.
 For instance, GPT-guided MCTS has enhanced search efficiency in symbolic regression [17], while value-guided MCTS decoding (PPO-MCTS [19]) improved the preferability of generated text over standard RLHF outputs.
 However, GPT-based MCTS approaches face notable pitfalls.
-Due to the autoregressive, token-by-token generation, even a single token error can derail an entire solution, and models often get stuck in suboptimal trajectories because their prior training biases lead to unbalanced exploration-exploitation.
+Due to autoregressive, token-by-token generation, even a single token error can derail an entire solution, and models often get stuck in suboptimal trajectories because their prior training biases lead to unbalanced exploration-exploitation.
 In other words, maintaining long-range coherence is challenging and it’s difficult to revise parts of a candidate without regenerating the whole sequence.
-Beyond these GPT-specific pitfalls, integrating MCTS into large-scale generative frameworks—particularly for biological design—faces major challenges as follows: First, unlike classical planning environments, we lack a tractable forward model to evaluate partially generated sequences.
-While recent approaches like Diffuser [9] show how diffusion models can operate as planners without explicit forward simulators, standard diffusion decoding is still largely open-loop.
-This limits its ability to adaptively revise candidates during generation.
-Second, real-world biological design tasks are intrinsically multi-objective.
-Protein design, for instance, requires satisfying multiple criteria, from structural compatibility to folding efficiency and evolutionary plausibility.
-Third, GPT-based generators make it inefficient to apply MCTS, especially for long-sequence tasks, resulting in unnecessarily large tree searches in both breadth and depth and inevitably exploring unpromising states.
-Finally, relying on a single GPT for exploration limits biophysical-fidelity, diversity, and constrains the search space due to biases from the pretraining dataset.
+Beyond these GPT-specific pitfalls, applying MCTS to biological sequence design raises core problems: (i) no reliable partial-state evaluator (i.e., no tractable forward model) to score or prune partial sequences; (ii) multi-objective constraints (e.g., structural compatibility, folding efficiency, evolutionary plausibility); (iii) poor scaling to long sequences under autoregressive generation, making tree search expensive and prone to exploring unpromising states; and (iv) limited diversity/biophysical fidelity when exploration relies on a single pretrained model.
+While Diffuser [11] shows diffusion models can act as planners without explicit forward simulators, diffusion decoding is still largely open-loop, offering limited opportunities for test-time correction.
 In this paper, we propose Monte Carlo Tree Diffusion with Multiple Experts (MCTD-ME), a novel planning framework that integrates diffusion-based generation with an MCTS search guided by a diverse ensemble of domain experts models.
@@
 Disagreement among experts encourages exploration; consensus promotes exploitation—enabling informed multi-objective planning at test time.
-We defer additional details on related work to Appendix 2.
-Summary of contributions: To our knowledge, this is the first work to incorporate an ensemble of expert into a diffusion-based planning framework, enabling multi-experts Monte Carlo tree search for complex generative problems.
-By unifying diffusion models with a multi-critic evaluation module and a multi-expert proposal module within MCTS, we introduce a general test-time strategy to steer generation toward goals that single decoders or static samplers struggle with—especially in large, structured domains like protein sequence design, where generation spaces are vast and require both parallel decoding and multi-perspective exploration.
-Central to our design are (i) a pLDDT-guided masking policy that focuses edits on structurally uncertain regions to improve biophysical fidelity, and (ii) a multi-expert selection mechanism that arbitrates among experts using cached uncertainty statistics, yielding stable and efficient search.
+Summary of contributions: To our knowledge, this is the first work to incorporate an ensemble of experts into a diffusion-based planning framework, enabling multi-expert Monte Carlo tree search for complex generative problems.
+By unifying diffusion models with a multi-critic evaluation module and a multi-expert proposal module within MCTS, we introduce a general inference-time strategy to steer generation toward goals that single decoders or static samplers struggle with—especially in large, structured domains like protein sequence design, where generation spaces are vast and require both parallel decoding and multi-perspective biophysical-fidelity-aware exploration.
 Second, we propose several technical innovations to realize this integration:
   1. 1.
-    A tree-structured diffusion process that expands and evaluates partial sequences through iterative biophysical-fidelity-enhanced denoising, achieved by masking low-pLDDT positions and guided by a multi-expert rollout mechanism, where each expert proposes candidate completions conditioned on the partial sequence, thereby enabling more diverse and meaningful exploration across a vast search space.
+    A tree-structured diffusion process for partial sequences, where each node represents a masked-infill state and expansion corresponds to a denoising rollout.
+We improve biophysical fidelity via a pLDDT-guided masking policy that focuses edits on structurally uncertain regions, and we enable diverse, goal-directed exploration using a multi-expert rollout mechanism, where each expert proposes candidate completions conditioned on the partial sequence.
   2. 2.
-    A novel $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$ selection algorithm, an entropy-aware selection rule, extending the classic UCT formula with exploration bonuses based on expert disagreement (e.g., score variance or entropy), inspired by entropy-guided UCB from ERP [21].
-Third, we demonstrate empirical gains on long-horizon generation tasks—specifically inverse protein folding for just one case example.
-Using the CAMEO 2022 benchmark [10] and a PDB date-split benchmark [4], we show that our framework consistently improves sequence recovery (AAR) and structural similarity (scTM) over strong DPLM-2 baselines [31].
-Our approach yields higher-quality, more diverse solutions, and these improvements grow with larger search budgets—highlighting the scalability and deliberative nature of our framework.
-We observe that both pLDDT-guided masking and multi-expert selection contribute complementary gains, with larger benefits on longer proteins where uncertainty is spatially concentrated.
+    A novel $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$ selection rule for multi-expert MCTS, which arbitrates among experts using cached uncertainty statistics for stable and efficient search, extending the classic UCT formula with exploration bonuses based on expert disagreement (e.g., score variance or entropy), inspired by entropy-guided UCB [23].
+Third, we demonstrate empirical gains on long-horizon generation tasks—folding, inverse folding, and motif scaffolding—on CAMEO 2022 [12] and a PDB date-split benchmark [5].
+In the lead-optimization regime (equal test-time compute), our planner consistently improves over strong baselines such as ESMFold [18], DPLM-2 [37], and ProteinMPNN [7] (we report absolute baseline/final values and $\Delta$).
+We obtain lower RMSD and higher TMscore in folding, higher AAR and scTM in inverse folding, and improved motif preservation in scaffolding; moreover, we outperform single-expert ablations of RFDiffusion [39], FoldFlow [4], and ProteinA [9], with gains amplifying on longer proteins and more challenging scaffolds where uncertainty is concentrated.
+Beyond lead optimization, we additionally evaluate a de novo design setting (all-mask root) in Appendix A.8.
+Our results highlight three practical advantages: (i) pLDDT-guided masking concentrates compute on low-confidence regions, improving efficiency and stability; (ii) multi-expert selection leverages complementary pretrained models, yielding robust gains even when individual experts underperform; and (iii) quality improves smoothly with larger search budgets, enabling controllable compute–performance trade-offs.
+Importantly, the method is model-agnostic and modular: it can integrate any set of pretrained experts and black-box critics at test time without retraining, making it broadly applicable across domains.
 Overall, this work advances the state-of-the-art in test-time guided generative planning, opening up promising directions for applications in drug discovery, program synthesis, and beyond.
 2 Related works
-We review related work in protein diffusion models, MCTS in generative modeling and molecule design, discrete diffusion for planning and generation, and multi-objective and expert-guided Generation.
-Protein and Drug Diffusion Models
-Recent advances apply diffusion models to protein design across both structure and sequence domains.
-RFdiffusion [32] introduced a guided 3D diffusion model for protein backbones, achieving state-of-the-art results.
-On the sequence side, DPLM [30] and its extension DPLM-2 [31] demonstrated that discrete diffusion pretraining on large-scale protein data enables high-quality sequence generation and versatile conditioning (e.g., inverse folding, property steering).
-Most recently, DiffGui [8] proposed a target-aware, E(3)-equivariant diffusion framework for joint atom and bond generation, incorporating multi-property guidance (e.g., binding affinity, drug-likeness) to generate realistic, high-affinity 3D molecules.
-However, these diffusion approaches generally rely on one-shot or open-loop decoding, making it difficult to incorporate adaptive search or revision during generation—highlighting the need for planning methods such as MCTS.
-Monte Carlo Tree Search in Generative Modeling
-Monte Carlo Tree Search (MCTS) has improved generative decoding in tasks requiring complex reasoning or program synthesis.
-In code generation, AlphaCode [14] used a brute-force generate-and-test method by sampling large program sets and selecting those that passed tests.
-More structured methods followed: RethinkMCTS [13] performs MCTS over LLM reasoning steps, using execution feedback to refine intermediate thoughts.
-Tree-of-Thoughts [33] frames reasoning as a tree of self-evaluated partial solutions expanded via classical search (DFS/BFS), boosting performance on puzzles and planning.
-PG-TD [34] simulates lookahead in code generation by executing candidate programs during decoding.
-Across domains, MCTS-style planning consistently enhances generation quality and reliability.
-Yet, most of these approaches are tied to autoregressive GPT-style models, which remain token-by-token and struggle with long-range dependencies—suggesting that discrete diffusion could provide a more flexible backbone for planning.
-Discrete Diffusion for Planning and Generation
-While originally developed for continuous domains, diffusion models have been successfully adapted to discrete sequence generation and planning.
-D3PM [1] introduced absorbing-state (masking) noise for discrete diffusion, achieving competitive text generation.
-Follow-ups like [27] further closed the performance gap with autoregressive models.
-In planning, Diffuser [9] modeled offline RL as trajectory denoising, generating full state-action paths guided by value functions.
-To enable adaptive inference, Monte Carlo Tree Diffusion (MCTD) [2] reimagines diffusion as tree search, branching multiple denoising outcomes per step and scoring partial trajectories via heuristics.
-In protein design, ProtInvTree [17] similarly integrates reward-guided search with jumpy denoising for efficient inverse folding.
-Still, existing methods optimize toward a single diffusion generator, which limits their ability to effective search space exploration in biological and chemical design.
-Multi-Experts Guided Generation
-Multi-expert systems leverage several generative models or “experts” in tandem, allowing each to contribute from its area of strength.
-This idea has been explored in reinforcement learning and bandit settings, where algorithms learn from multiple expert policies or oracles to improve decision-making [20].
-In generative modeling, mixture-of-experts strategies similarly combine specialized models to better cover complex data distributions.
-For example, recent diffusion-based image generators use multiple specialized denoising models across different noise levels [5] or even heterogeneous model architectures [12] to boost output quality.
-However, these multi-expert approaches remain largely static and one-shot during inference, without an adaptive search process.
-To our knowledge, no prior diffusion model integrates multiple experts into a planning loop for generative design.
-By incorporating a team of expert generators within an MCTS-driven diffusion framework, our work enables dynamic collaboration among experts to satisfy multiple design criteria and explore a broader solution space—an innovative combination that extends diffusion-based generation to more complex, multi-objective scenarios.
+Recent work shows diffusion models are promising for biomolecular design, including structure-based backbone generation (RFdiffusion [38]), discrete sequence generation with versatile conditioning (DPLM and DPLM-2 [36, 37]), and property-guided atom–bond generation (DiffGui [10]); however, these methods are typically one-shot and lack test-time refinement.
+In parallel, tree search has improved generation for code and reasoning (AlphaCode [16], RethinkMCTS [15], Tree-of-Thoughts [40], PG-TD [43]) and more recently System-2-style planning for LMs [2], but largely targets autoregressive GPT-style models rather than diffusion.
+Discrete diffusion supports generation via absorbing-state noise (D3PM [1]) and masked-diffusion LMs simplifying denoising [32, 31], plus planning (Diffuser [11]), inverse folding (ProtInvTree [20]), and peptide optimization (PepTune [34]).
+Closest, masked-diffusion has also been used for planning in diffusion spaces (e.g., path planning) [31, 32], but these methods plan over trajectories, whereas our setting is protein sequence design with biophysical critics and pLDDT-guided edit control.
+While mixture-of-experts strategies have been explored in RL and diffusion-based image generation [8, 14, 22], they are static at inference.
+Our approach integrates multiple specialized expert generators within an MCTS-driven diffusion framework for dynamic collaboration, adaptive search, and multi-objective biomolecular design.
+(Additional details are provided in Appendix A.6.)
 3 Preliminaries
-We describe Diffusion-based generative model, MCTS, and protein design along with their mathematical notation in the subsequent sections and unify them under the Markov decision processes framework.
 Markov decision processes.
-We formulate generative planning as a finite-horizon MDP $\mathcal{M}=\langle\mathcal{S},\mathcal{A},H,P,R\rangle$ [26], where $\mathcal{S}$ is the set of states (e.g., partial or completed sequences), $\mathcal{A}$ is the action space (e.g., denoising steps or token insertions), and $P:\mathcal{S}\times\mathcal{A}\to\mathcal{S}$ defines stochastic transitions.
-The reward function $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ provides a scalar evaluation, often defined only on terminal states.
-The policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ defines the action distribution at each state, such as a denoising model in diffusion or a next-token model in autoregressive generation.
-An episode corresponds to constructing a sequence $y=(y_{1},...,y_{H})$, with a termination action marking the end of generation.
-We denote the action-value function $Q^{\pi}(s,a)$ as the expected cumulative reward of taking action $a$ in state $s$ under policy $\pi$, and the value function $V^{\pi}(s)$ as the expected return from $s$.
+We formulate generative planning as a finite-horizon MDP $\mathcal{MDP}=\langle\mathcal{S},\mathcal{A},H,P,R\rangle$ [30], where $\mathcal{S}$ is the set of states (e.g., partial or completed sequences), $\mathcal{A}$ is the action space (e.g., denoising steps or token insertions), and $P:\mathcal{S}\times\mathcal{A}\to\mathcal{S}$ defines stochastic transitions.
+The reward $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is often defined only on terminal states, and the policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ specifies an action distribution (e.g., a diffusion denoiser or autoregressive next-token model).
+An episode constructs a length-$H$ sequence $y=(y_{1},...,y_{H})$.
+We denote the action-value function $Q^{\pi}(s,a)$ and value function $V^{\pi}(s)$.
 Diffusion-based generative model.
-We briefly review the diffusion framework for sequence generation.
-Let $y_{0:H}$ denote a token sequence of length $H$.
-A forward noising process $q(y_{t}\mid y_{t-1})$, for $t=1,\ldots,T$, gradually corrupts the sequence, such that $y_{T}$ becomes maximally noisy (e.g., fully masked for discrete data or Gaussian noise for continuous).
-A reverse denoising model $p_{\phi}(y_{t-1}\mid y_{t})$ is trained to recover $y_{t-1}$ from $y_{t}$, enabling sampling via iterative denoising from $y_{T}\sim p(y_{T})$ down to $y_{0}\sim p_{\phi}(y_{0})$.
-Crucially, diffusion models naturally support guided generation.
+We review diffusion for sequence generation.
+Let $y_{0:H}$ denote an $H$-token sequence.
+Forward noising $q(y_{t}\mid y_{t-1})$ for $t=1,\ldots,T$ corrupts the sequence until $y_{T}$ is maximally noisy (e.g., fully masked).
+A reverse model $p_{\phi}(y_{t-1}\mid y_{t})$ is trained to denoise, enabling sampling by iterating from $y_{T}\sim p(y_{T})$ to $y_{0}\sim p_{\phi}(y_{0})$.Crucially, diffusion models naturally support guided generation.
 When an auxiliary score function $f(y)$ is available—e.g., a property predictor or constraint critic—generation can be biased toward desirable outputs.
@@
 $p^{\text{guide}}_{\phi}(y_{t-1}\mid y_{t})\propto p_{\phi}(y_{t-1}\mid y_{t})\cdot\exp\left\{\beta\cdot f(y_{t-1})\right\},$
-with $\beta$ controlling guidance strength.
-This formulation generalizes to multiple critics: given $K$ evaluators $C_{0},\dots,C_{K-1}$, each scoring complete sequences, guidance becomes even more natural.
-Instead of aggregating into a scalar reward, we perform imitation-style rollouts: partial candidates are completed, scored individually by each expert, and used to inform node evaluations during tree search.
-This preserves the diversity of expert preferences and supports structured planning, which we realize through our Monte Carlo Tree Diffusion (MCTD) framework.
+with $\beta$ controlling guidance strength. We extend this to multiple experts: $K$ evaluators $\{C_{k}\}_{k=1}^{K}$ score complete sequences proposed by multiple experts $\{\pi^{e}\}_{e=1}^{E}$ .
 MCTS.
-Monte Carlo Tree Search (MCTS) is a planning algorithm that combines tree-based exploration with stochastic simulation to balance exploration and exploitation in large decision spaces.
-It proceeds in four stages: selection, expansion, simulation, and backpropagation.
-Starting from the root (current state), the selection phase recursively chooses child nodes using an Upper Confidence Bound (UCT) criterion [11]:
+Monte Carlo Tree Search (MCTS) balances exploration and exploitation via tree search and simulation. It iterates selection, expansion, evaluation, and backpropagation. Selection uses UCT [13]:
 $\text{UCB}=Q\left({s,a}\right)+c_{p}\cdot\sqrt{\frac{\log\left({N\left({s}\right)}\right)}{N\left({s,a}\right)}},$ (1)
 where $Q(s,a)$ is the estimated reward, $N(s)$ and $N(s,a)$ are visit counts, and $c_{p}$ controls exploration.
-When a leaf node is reached, expansion adds child nodes corresponding to unexplored actions.
-Then, simulation evaluates these new nodes by executing a rollout policy (e.g., sampling until termination or a predefined depth), producing a scalar reward.
-Finally, backpropagation propagates the reward up the visited path, updating $N(\cdot)$ and $Q(\cdot)$ statistics.
-Over repeated iterations, MCTS refines its value estimates and focuses the search on high-reward regions of the state space.
+Expansion adds children, evaluation runs a rollout to obtain a reward, and backpropagation updates $N(\cdot)$ and $Q(\cdot)$.
+Repetition concentrates search on high-reward regions.
 4 The MCTD-ME Algorithm
-Figure 1: Overview of MCTD in diffusion-based protein sequence generation. Nodes are partially denoised sequences expanded via masked diffusion and guided by multi-expert rollouts.
+Figure 1: Overview of MCTD-ME in diffusion-based protein sequence generation. Nodes are partially denoised sequences expanded via masked diffusion and guided by multi-expert rollouts.
 We propose Monte Carlo Tree Diffusion with Multi-Experts (MCTD-ME), which enhances protein generation by combining masked diffusion with an MCTS planner.
+An overview of the MCTD-ME workflow is shown in Fig 1.
 MCTD-ME structures diffusion as a tree, refining partially masked sequences to preserve scaffold integrity while systematically exploring uncertain regions.
-Search quality and diversity are further improved by incorporating multiple diffusion experts for diverse refinements and by using pLDDT-guided masking to target low-confidence residues.
+Search quality and diversity are further improved by incorporating multiple diffusion experts for diverse refinements and by using pLDDT-guided masking to target low-confidence residues, as illustrated in Fig 2.
 These design choices couple the planning strength of tree search with expert-aware diversification, yielding a principled framework for balancing exploration and exploitation in protein modeling.
-We present the algorithm in Algorithm 1, with each rollout consisting of four key steps.
+We present the algorithm in Algorithm 1 (deferred to Appendix A.1), with each rollout consisting of four key steps as following:
 Selection.
 To navigate the space of protein sequences under the reverse masked diffusion framework, we define a tree where each node corresponds to a complete sequence (with no masked tokens) at a particular reverse step, and edges represent one reverse transition obtained by applying masked discrete diffusion to a low-confidence subset of tokens.
-At each iteration, we employ a novel selection algorithm, $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$—an entropy-augmented variant of UCT—designed to identify the most promising path, where ME denotes Multiple Experts.
+At each iteration, we employ a novel selection algorithm, $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$—a Policy(P)-guided entropy($\mathcal{H}$)-augmented variant of UCT—designed to identify the most promising path, where ME denotes Multiple Experts.
 Specifically, from a parent node $s_{t}$ we choose the transition (i.e., reverse move to $s_{t-1}$ via a particular mask set) that maximizes the following objective:
-$\text{P}\mathcal{H}\text{-UCT}\text{-ME}(s_{t})\;=\;\mathop{\rm arg\,max}_{a\in\mathcal{A}}\;\text{P}\mathcal{H}\text{-UCB}\text{-ME}\!\left((s_{t},a)\right),$ (2)
+$\text{P}\mathcal{H}\text{-UCT}\text{-ME}(s_{t})\;=\;\mathop{\rm arg\,max}_{a\in\mathcal{A}}\;\text{P}\mathcal{H}\text{-UCB}\text{-ME}\!\left((s_{t},a)\right),$
 with
-$\text{P}\mathcal{H}\text{-UCB}\text{-ME}\!\left((s_{t},a)\right)$ $=Q(s_{t},a)+c_{p}\,\frac{\sqrt{\log N(s_{t})}}{1+N(s_{t},a)}\cdot\left({\underbrace{w_{\mathrm{ent}}\,\mathcal{U}_{\text{ent}}(s_{t},a)\;+\;w_{\mathrm{div}}\,\mathcal{U}_{\text{div}}(s_{t},a)}_{\text{$\text{P}\mathcal{H}$\text{-ME} bonus (cached at expansion)}}}\right).$ (3)
+$\text{P}\mathcal{H}\text{-UCB}\text{-ME}\!$ $\left((s_{t},a)\right)=Q(s_{t},a)+c_{p}\,\frac{\sqrt{\log N(s_{t})}}{1+N(s_{t},a)}$
+$\;\pi_{\mathrm{cons},\tau}(a\mid s_{t})$ $\cdot\left({\underbrace{w_{\mathrm{ent}}\,\mathcal{U}_{\text{ent}}(s_{t},a)\;+\;w_{\mathrm{div}}\,\mathcal{U}_{\text{div}}(s_{t},a)}_{\text{$\text{P}\mathcal{H}$\text{-ME}~bonus (cached at expansion)}}}\right).$ (2)
 Here, $Q(s_{t},a)$ is the current value estimate, $N(\cdot)$ are visit counts, $c_{p}>0$ controls exploration, and $w_{\mathrm{ent}},w_{\mathrm{div}}\!\geq\!0$ weight the predictive-uncertainty and diversity bonuses, respectively.
 In our diffusion setting, the action prior is given by the reverse diffusion model $p_{\phi}(s_{t-1}\!\mid\!s_{t})$ over masked edits, but—consistent with our implementation—we do not recompute its entropy during selection; instead we cache uncertainty statistics for each expanded child and reuse them during subsequent selections.
-Multi-expert PH bonus.
-Let $\mathcal{M}$ be the mask set chosen at $s_{t}$ (e.g., via pLDDT), and let $\{p_{\phi}^{(e)}\}_{e=1}^{E}$ denote the $E$ expert conditional distributions (softmax of logits) used during expansion.
-For the candidate child $s_{t-1}$ obtained by action $a$, we define the ensemble-based uncertainty as
-$\mathcal{U}_{\text{ent}}(s_{t},a)=\mathcal{H}\!\left(\frac{1}{E}\sum_{e=1}^{E}p_{\phi}^{(e)}(\,\cdot\,\mid s_{t-1},\mathcal{M})\right)-\frac{1}{E}\sum_{e=1}^{E}\,\mathcal{H}\!\left(p_{\phi}^{(e)}(\,\cdot\,\mid s_{t-1},\mathcal{M})\right).$ (4)
-Here $\mathcal{H}[p]=-\sum_{a}p(a)\log p(a)$ denotes Shannon entropy.
-The first term is the entropy of the ensemble-averaged distribution $\bar{p}=\tfrac{1}{E}\sum_{e}p^{(e)}$, while the second is the average entropy of the individual experts.
-Their difference equals the mutual information (BALD score [7]), or Jensen–Shannon–style divergence between experts.
-Intuitively, it isolates epistemic uncertainty: if experts disagree, $\mathcal{H}[\bar{p}]$ is large while the average $\mathcal{H}[p^{(e)}]$ remains small, giving a positive signal.
-If all experts are merely noisy but agree, the two terms cancel.
+We cache a temperature-controlled consensus prior $\pi_{\mathrm{cons},\tau}(a\mid s_{t})\!\in\![0,1]$ at expansion—large only when multiple experts assign high mass to the same action—so the UCB exploration is steered toward expert-agreed edits while preserving asymptotics.
+We define
+$\pi_{\mathrm{cons},\tau}(a\mid s)=\frac{1}{Z_{\tau}}\exp\!\Big(\tfrac{1}{\tau}\sum_{e=1}^{E}\log p^{(e)}_{\phi}(a\mid s,\mathcal{M})\Big),$ (3)
+$Z_{\tau}=\sum_{a^{\prime}\in\mathcal{A}(s)}\exp\!\Big(\tfrac{1}{\tau}\sum_{e=1}^{E}\log p^{(e)}_{\phi}(a^{\prime}\mid s,\mathcal{M})\Big),$ (4)
+i.e., a temperatured geometric mean across experts.
+[…]
+$\mathcal{U}_{\text{ent}}(s_{t},a)=$ $\mathcal{H}\!\left(\frac{1}{E}\sum_{e=1}^{E}p_{\phi}^{(e)}(\,\cdot\,\mid s_{t-1},\mathcal{M})\right)$
+$-\frac{1}{E}\sum_{e=1}^{E}\,\mathcal{H}\!\left(p_{\phi}^{(e)}(\,\cdot\,\mid s_{t-1},\mathcal{M})\right).$ (5)
+Here $\mathcal{H}(p)=-\sum_{a}p(a)\log p(a)$ is Shannon entropy; the difference equals a BALD/Jensen–Shannon–style disagreement signal, isolating epistemic uncertainty.
 We also include a diversity measure, defined as the normalized Hamming distance between parent and child sequences:
-$\mathcal{U}_{\text{div}}(s_{t},a)$ $=\frac{1}{L}\sum_{i=1}^{L}\,\mathbf{1}\{y^{\text{child}}_{i}\neq y^{\text{parent}}_{i}\},$ (5)
-For our inverse folding case studies, this simply counts the fraction of amino acids that differ.
-Both $\mathcal{U}_{\text{ent}}$ and $\mathcal{U}_{\text{div}}$ are computed once at expansion and stored for reuse during selection.
+$\mathcal{U}_{\text{div}}(s_{t},a)$ $=\frac{1}{L}\sum_{i=1}^{L}\,\mathbf{1}\{y^{\text{child}}_{i}\neq y^{\text{parent}}_{i}\}.$ (6)
+In our multi-model setting, sequences are batched and multi-channel (e.g., amino-acid tokens and structure tokens).
+We evaluate $\mathcal{U}_{\text{div}}$ over the task-relevant token subset: for folding, the difference is computed on structure tokens relative to the root structure; for inverse folding, on amino-acid tokens relative to the baseline sequence; and for motif scaffolding, on both channels (combined by concatenation or by averaging channel-wise normalized distances).
+When batching, we average the per-sequence $\mathcal{U}_{\text{div}}$ across the batch, thus the result remains in $[0,1]$.
+Both $\mathcal{U}_{\text{ent}}$ and $\mathcal{U}_{\text{div}}$ are computed once at expansion and cached for reuse during selection.
 An action $a$ is prioritized if it (i) has high estimated value $Q(s_{t},a)$, (ii) corresponds to an under-explored branch (UCB exploration), (iii) induces strong ensemble disagreement ($\mathcal{U}_{\text{ent}}$ large), indicating information gain, and (iv) introduces sufficient novelty relative to the parent ($\mathcal{U}_{\text{div}}$ large).
 This encourages the search to balance reward exploitation with uncertainty- and diversity-driven exploration, while the reverse diffusion prior $p_{\phi}(s_{t-1}\!\mid\!s_{t})$ maintains consistency with the denoising dynamics.
-Figure 2: Masked discrete diffusion step: low-confidence sites (mask) are resampled while confident tokens remain fixed, yielding a progressively “unmasked” sequence.
+Remark 1 (single-expert case).
+When E=1E{=}1, (5) vanishes ( $\mathcal{U}_{\text{ent}}{=}0$ ).
+In MCTD-1, we therefore use the predictive (Shannon) entropy of the single expert over the masked sites as the uncertainty bonus:
+$\mathcal{U}_{\text{ent}}^{(1)}(s_{t},a)=\frac{1}{|\mathcal{M}|}\sum_{i\in\mathcal{M}}\mathcal{H}\!\big(p_{\phi}(x_{i}\mid s_{t-1},\mathcal{M})\big),$
+which preserves the same units and interpretation (larger $\Rightarrow\,$ more uncertain edit).
+This value is computed at expansion and cached for reuse, akin to the entropy-guided UCB bonus used in ERP [23].
+In the single-expert case, the multiplicative prior reduces to that expert’s $\pi_{\tau}(a\mid s_{t})$, ensuring a coherent reduction.
 Expansion.
@@
 Rather than resampling all positions uniformly, we exploit predicted structural confidence (pLDDT) to target only low-confidence residues.
-Specifically, at reverse step $t$ we define a mask $\mathcal{M}_{t}$ by thresholding the pLDDT profile of $y_{t}$, masking unstable regions while preserving high-confidence subsequences.
+Specifically, at reverse step $t$ we define a mask $\mathcal{M}t$ by thresholding the pLDDT profile of $y_{t}$, masking unstable regions while preserving high-confidence subsequences.
 Crucially, the masking threshold decreases over diffusion steps, progressively reducing the fraction of masked residues as denoising proceeds.
@@
 $y_{t-1}\sim p_{\phi}(y_{t-1}\mid\text{Mask}(y_{t};\mathcal{M}_{t})).$ (7)
-This masked input encourages structural refinement and prevents overwriting already stable subsequences, reducing unnecessary sampling noise and improving convergence.
+Mixture-over-experts view. Conceptually, expansion implements a guided sampler that averages proposal distributions from $E$ pretrained experts and biases them by reward $R$:
+$p^{\text{plan}}(y_{t-1}\mid y_{t})\ \propto\$$\Big[\sum_{e=1}^{E}w_{e}p^{(e)}\phi(y_{t-1}\mid y_{t},\mathcal{M}_{t})\Big]$
+$\cdot\exp\{\beta,R(y_{t-1})\},$ (8)
+where $p^{(e)}_{\phi}$ is expert $e$’s conditional, $w_{e}$ are mixing weights, and $\beta$ controls guidance.
+In practice we realize this by sample–score–select (below) rather than gradient injection.
 Multi-Expert Guided Expansion via Rollouts.
-To enrich proposal diversity and better imitate the training distribution, we leverage multiple pretrained experts during expansion.
-Given a masked sequence $\text{Mask}(y_{t};\mathcal{M}_{t})$, each expert $e\in\{1,\dots,E\}$ performs $k$ rollouts by sampling its conditional distribution $p_{\phi}^{(e)}(\cdot\mid y_{t},\mathcal{M}_{t})$.
+To enrich proposal diversity and imitate the training distribution, we leverage multiple experts during expansion.
+Given a masked sequence $\text{Mask}(y_{t};\mathcal{M}_{t})$, each expert $\pi^{e}\in\{1,\dots,E\}$ performs $k$ rollouts by sampling its conditional distribution $p_{\phi}^{(e)}(\cdot\mid y_{t},\mathcal{M}_{t})$.
 This yields a pooled candidate set
 $\mathcal{C}_{t}\;=\;\bigcup_{e=1}^{E}\ \bigcup_{r=1}^{k}\ y_{t-1}^{(e,r)},$ (9)
-where $y_{t-1}^{(e,r)}$ denotes a complete sequence obtained by filling the masked sites and splicing them back into $y_{t}$.
-All candidates in $\mathcal{C}_{t}$ are de-duplicated and scored once with the composite reward (see Evaluation).
-We then retain the top-$K$ children, each annotated with two exploration bonuses computed at expansion: an ensemble surprisal term $\mathcal{U}_{\text{ent}}$ (capturing expert disagreement on the masked sites) and a novelty term $\mathcal{U}_{\text{div}}$ (the normalized Hamming distance to the parent).
-These cached quantities are reused during $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$  selection, ensuring that search is guided by both expert-informed proposal diversity and explicit uncertainty signals.
+where $y_{t-1}^{(e,r)}$ is formed by filling masked sites and splicing back into $y_{t}$.
+We de-duplicate $\mathcal{C}_{t}$, compute a composite score $R(\cdot)$ (see Evaluation) per candidate, cache the results, and retain the top-$K$ children by $R$.
+Each kept child stores two exploration bonuses computed once at expansion—ensemble surprisal $\mathcal{U}_{\text{ent}}$ and novelty $\mathcal{U}_{\text{div}}$—which are reused during $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$ selection.
+Figure 2: Masked discrete diffusion step: low-confidence sites (mask) are resampled while confident tokens remain fixed, yielding a progressively “unmasked” sequence.
 Candidate Proposal Strategy. To retain diversity while prioritizing promising edits, we rank all expanded children and select the top-$K$ candidates:
 $\left\{y_{i=1}^{K}\right\}=\operatorname{Top\text{-}K}\big(\{y_{t-1}^{(j)}\},K\big).$ (10)
-Here, $j$ indexes every candidate child pooled across experts and rollouts at step $t$; $\operatorname{Top\text{-}K}$ ranks them by the composite score and returns the best $K$.
-Each chosen child is inserted into the tree as a new node, annotated with its cached $\text{P}\mathcal{H}\text{-ME}$ bonuses ($\mathcal{U}_{\text{ent}},\mathcal{U}_{\text{div}}$) and expert rollout scores.
+Here, $j$ indexes all candidates pooled across experts and rollouts at step $t$; $\operatorname{Top\text{-}K}$ ranks by the composite score $R$ and returns the best $K$.
+Each chosen child is inserted into the tree as a new node, annotated with its cached $\text{P}\mathcal{H}\text{-ME}$ bonuses ($\mathcal{U}_{\text{ent}},\mathcal{U}_{\text{div}}$) and expert rollout score.
 This beam-style expansion ensures that multiple high-quality hypotheses are pursued in parallel, enabling MCTD-ME to balance exploitation of strong candidates with exploration of diverse structural alternatives.
+Configuration RMSD $\downarrow$ TM-score $\uparrow$ Composite $\uparrow$
+Base Final $\Delta$ Base Final $\Delta$ Base Final $\Delta$
+CAMEO (Folding evaluation)
+Baseline One-shot (150M) $10.50\pm 6.42$ $0.537\pm 0.184$ $0.302\pm 0.228$
+Random (MCTS-0) $10.53\pm 6.40$ $0.14\pm 2.66$ $0.813\pm 0.058$ $0.281\pm 0.153$ $0.403\pm 0.253$ $0.107\pm 0.132$
+Single-Expert (150M) $10.00\pm 6.62$ $0.50\pm 1.65$ $0.809\pm 0.060$ $0.272\pm 0.150$ $0.426\pm 0.253$ $0.124\pm 0.099$
+Single-Expert (650M) $9.24\pm 6.11$ $1.29\pm 2.02$ $0.811\pm 0.061$ $0.273\pm 0.148$ $0.450\pm 0.248$ $0.147\pm 0.103$
+Single-Expert (3B) $9.73\pm 6.49$ $0.79\pm 2.07$ $0.810\pm 0.061$ $0.271\pm 0.149$ $0.434\pm 0.257$ $0.132\pm 0.108$
+Sampling $9.97\pm 6.43$ $0.52\pm 1.58$ $0.804\pm 0.064$ $0.264\pm 0.146$ $0.422\pm 0.252$ $0.118\pm 0.093$
+MCTD-UCT $9.47\pm 6.64$ $1.28\pm 2.23$ $0.827\pm 0.063$ $0.296\pm 0.151$ $0.454\pm 0.262$ $0.160\pm 0.113$
+MCTD-ME $\mathbf{9.41\pm 6.65}$ $\mathbf{+1.32\pm 2.33}$ $\mathbf{0.827\pm 0.062}$ $\mathbf{+0.296\pm 0.154}$ $\mathbf{0.456\pm 0.263}$ $\mathbf{+0.160\pm 0.119}$
+PDB (Folding evaluation)
+Baseline One-shot (150M) $8.45\pm 5.41$ $0.591\pm 0.177$ $0.384\pm 0.233$
+Random (MCTS-0) $8.07\pm 5.12$ $0.37\pm 2.29$ $0.705\pm 0.043$ $0.114\pm 0.190$ $419\pm 0.176 +0.0$$035\pm 0.124$
+Single-Expert (150M) $7.55\pm 5.35$ $0.90\pm 1.94$ $0.705\pm 0.042$ $0.114\pm 0.186$ $439\pm 0.185 +0.0$$055\pm 0.116$
+Single-Expert (650M) $6.84\pm 5.08$ $1.66\pm 2.48$ $0.705\pm 0.042$ $0.114\pm 0.185$ $464\pm 0.176 +0.0$$081\pm 0.136$
+Single-Expert (3B) $7.27\pm 5.41$ $1.24\pm 2.37$ $0.706\pm 0.042$ $0.114\pm 0.186$ $450\pm 0.187 +0.0$$067\pm 0.130$
+Sampling $7.63\pm 5.27$ $0.82\pm 1.82$ $0.703\pm 0.039$ $0.110\pm 0.185$ $435\pm 0.182 +0.0$$050\pm 0.114$
+MCTD-UCT $6.73\pm 5.18$ $1.68\pm 2.44$ $0.711\pm 0.047$ $0.117\pm 0.189$ $471\pm 0.180 +0.0$$085\pm 0.135$
+MCTD-ME $\mathbf{6.47\pm 5.22}$ $\mathbf{+1.97\pm 2.52}$ $\mathbf{0.720\pm 0.047}$ $\mathbf{+0.129\pm 0.188}$ $\mathbf{0.471\pm 0.180}$ $\mathbf{+0.084\pm 0.137}$
+[…]
 Evaluation.
-Unlike expansion, where multiple experts serve as proposal generators, the evaluation stage relies on a set of critics to score complete sequences.
-Because every node in the tree corresponds to a fully denoised sequence $y_{0}$, thus evaluation can be performed immediately without further decoding.
-Each sequence is scored by a panel of critics $\{C_{1},\dots,C_{J}\}$, producing values
-$C_{j}(y_{0})\;:=\;R_{j}(y_{0}),\quad j=1,\dots,J,$ (9)
-where the $C_{j}$ represent evaluation functions such as sequence recovery (AAR), structural similarity (scTM), or biophysical quality metrics, $J$ here denotes the number of critics.
-In practice, we aggregate these scores into a single scalar via a fixed weighted sum
-$R(y_{0})\;=\;\sum_{j=1}^{J}w_{j}\,C_{j}(y_{0}),\qquad w_{j}\geq 0,\ \sum_{j=1}^{J}w_{j}=1,$ (10)
-This design matches the diffusion context, where evaluation is only meaningful at complete denoised sequences.
+Whereas expansion uses experts to generate proposals, evaluation uses a panel of critics to score complete sequences.
+Because every node in the tree corresponds to a fully denoised sequence $y_{0}$, evaluation can be performed immediately without further decoding.
+Each sequence is scored by critics $C_{1},\dots,C_{J}$}, producing
+$C_{j}(y_{0})\coloneq R_{j}(y_{0}),\quad j=1,\dots,J.$ (11)
+We normalize critic outputs to $[0,1]$ and aggregate via a fixed convex combination
+$R(y_{0})=\sum_{j=1}^{J}w_{j},C_{j}(y_{0}),\qquad w_{j}\geq 0,\ \sum_{j=1}^{J}w_{j}=1,$ (12)
+then cache $R(y_{0})$ for re-use in selection/backprop.
+This separation—experts for proposals, critics for scores—implements a tractable, planning-based approximation to the guided reverse process sketched above.
 Backpropagation.
@@
 This max-based backup, consistent with our implementation, favors exploiting high-reward branches while maintaining the ability to explore alternative completions.
-Because each node is a complete sequence, long-range credit assignment is naturally aligned with the denoising trajectory.
-Variant AAR Normalized Reward scTM
-CAMEO Benchmark
+Because each node is a complete sequence, long-range credit assignment is aligned with the denoising trajectory.
+[…]
+5 Experiments
+Variant AAR $\uparrow$ Norm. Reward $\uparrow$ scTM $\uparrow$
 Baseline Final $\Delta$ Baseline Final $\Delta$ Baseline Final $\Delta$
-MCTD-ME-0 (Random) 0.438 0.381 0.057-0.057 0.425 0.429 +0.004+0.004 0.362 0.458 +0.096+0.096
-Single-Expert (150M, Exp. 1) 0.427 0.427 +0.000+0.000 0.290 0.290 +0.000+0.000 0.371 0.371 +0.000+0.000
-Single-Expert (650M, Exp. 0) 0.442 0.458 +0.016+0.016 0.431 0.447 +0.016+0.016 0.372 0.387 +0.014+0.014
-Single-Expert (3B, Exp. 2) 0.443 0.429 0.014-0.014 0.432 0.422 0.011-0.011 0.373 0.369 0.005-0.005
-MCTD-ME (Multi-Expert) 0.439 0.451 +0.012 0.427 0.465 +0.038 0.365 0.449 +0.084
-PDB Benchmark
-MCTD-ME-0 (Random) 0.430 0.400 0.030-0.030 0.110 0.100 0.010-0.010 0.360 0.380 +0.020+0.020
-MCTD-ME-1 (650M) 0.490 0.520 +0.030+0.030 0.330 0.360 +0.030+0.030 0.390 0.420 +0.030+0.030
-MCTD-ME-1 (150M) 0.460 0.460 +0.000+0.000 0.290 0.290 +0.000+0.000 0.370 0.370 +0.000+0.000
-MCTD-ME-1 (3B) 0.500 0.420 0.080-0.080 0.340 0.280 0.060-0.060 0.400 0.370 0.030-0.030
-MCTD-ME (Multi-Expert) 0.500 0.560 +0.060 0.350 0.410 +0.060 0.880 0.910 +0.030
-Table 1: Main results on CAMEO and PDB.
-Baseline and final values for AAR, normalized reward, and scTM are shown, with $\Delta$ indicating improvements.
-Multi-expert MCTD consistently outperforms single-expert and unguided baselines, especially on PDB targets.
-5 Experiments
+CAMEO
+DPLM2-150M (Single-Expert, Baseline) $0.4425\pm 0.1190$ $0.3893\pm 0.1326$ $0.3701\pm 0.1528$
+Random (MCTD-0) $4250\pm 0.009 0.0$$0175\pm 0.0145$ $3726\pm 0.008 0.0$$0167\pm 0.0130$ $3763\pm 0.014 +0.0$$0062\pm 0.0560$
+DPLM2-650M (Single-Expert) $4535\pm 0.1162 +0.0$$0110\pm 0.0149$ $3974\pm 0.1010 +0.0$$0081\pm 0.0148$ $3621\pm 0.1515 0.0$$0080\pm 0.0678$
+ProteinMPNN (Single-Expert) $4326\pm 0.1102 0.0$$0098\pm 0.0194$ $3786\pm 0.1026 0.0$$0107\pm 0.0304$ $3618\pm 0.1468 0.0$$0070\pm 0.0691$
+Sampling (Multi-Expert) $4264\pm 0.1281 0.0$$0161\pm 0.0511$ $4113\pm 0.0896 +0.0$$0220\pm 0.0376$ $4194\pm 0.1536 +0.0$$0593\pm 0.0729$
+MCTD-UCT (Multi-Expert) $4465\pm 0.1170 +0.0$$0040\pm 0.0397$ $4129\pm 0.0851 +0.0$$0236\pm 0.0307$ $3895\pm 0.1545 +0.0$$0098\pm 0.0745$
+MCTD-ME (Multi-Expert) $\mathbf{0.4667\pm 0.1124}$ $\mathbf{+0.0237\pm 0.0238}$ $\mathbf{0.4132\pm 0.1119}$ $\mathbf{+0.0239\pm 0.0461}$ $\mathbf{0.4204\pm 0.1585}$ $\mathbf{+0.0503\pm 0.0505}$
+PDB
+DPLM2-150M (Single-Expert, Baseline) $0.5037\pm 0.1196$ $0.4629\pm 0.0763$ $0.3804\pm 0.1575$
+Random (MCTD-0) $4876\pm 0.1119 0.0$$0160\pm 0.0154$ $4591\pm 0.0751 0.0$$0032\pm 0.0157$ $3773\pm 0.1585 0.0$$0030\pm 0.0505$
+DPLM2-650M (Single-Expert) $5130\pm 0.1208 +0.0$$0109\pm 0.0141$ $4751\pm 0.0812 +0.0$$0122\pm 0.0134$ $3817\pm 0.1594 +0.0$$0022\pm 0.0490$
+ProteinMPNN (Single-Expert) $4882\pm 0.1112 0.0$$0155\pm 0.0163$ $4594\pm 0.0748 0.0$$0029\pm 0.0143$ $3774\pm 0.1575 0.0$$0030\pm 0.0596$
+Sampling (Multi-Expert) $4927\pm 0.1982 0.0$$0110\pm 0.0289$ $4790\pm 0.0728 +0.0$$0161\pm 0.0140$ $3844\pm 0.1536 +0.0$$0040\pm 0.0527$
+MCTD-UCT (Multi-Expert) $5219\pm 0.1388 +0.0$$0182\pm 0.0189$ $4793\pm 0.0881 +0.0$$0164\pm 0.0198$ $3816\pm 0.1545 +0.0$$0012\pm 0.0536$
+MCTD-ME (Multi-Expert) $\mathbf{0.5244\pm 0.1350}$ $\mathbf{+0.0213\pm 0.0139}$ $\mathbf{0.4828\pm 0.0900}$ $\mathbf{+0.0199\pm 0.0146}$ $\mathbf{0.3827\pm 0.1549}$ $\mathbf{+0.0033\pm 0.0410}$
+Table 2: Inverse folding on CAMEO and PDB with a fixed lead.
+For each target, the baseline lead is the one-shot sequence from DPLM-2 (150M) (first row in each dataset block); Final is the sequence after refinement, and $\Delta$ is computed per target against the same fixed baseline, isolating optimization effects from lead-sampling noise.
+Single-Expert denotes refinement using a single proposal generator (e.g., DPLM-2 650M or ProteinMPNN) rather than an ensemble.
+MCTD-0 uses random expert routing.
+Sampling (Multi-Expert) draws proposals from all experts on the pLDDT-derived mask without tree search (no selection/backpropagation), and MCTD-UCT (Multi-Expert) runs MCTS with standard UCB selection but without the uncertainty/novelty bonuses used by MCTD-ME.
+MCTD-ME (Multi-Expert) is our full method with informed expert selection and uncertainty-aware edits.
+Metrics are amino acid recovery (AAR), normalized reward, and scTM (higher is better).
+Values are mean $\pm$ s.e.m.
+across targets (CAMEO n=183n{=}183, PDB n=449n{=}449).
+Bold indicates the best within each dataset block.
+5.1 Experimental Configuration
+Models, experts, and baselines.
+We use masked discrete diffusion rollouts and instantiate MCTD-ME with task-specific proposal experts: sequence experts are pre-trained DPLM-2 models [36] of different capacities (150M, 650M, 3B), and motif-scaffolding experts include structure-aware models (RFDiffusion [39], ProteinA [9], FoldFlow [4]).
+All experts propose from the same masked input and are scored by a single composite critic.
+We evaluate a lead-optimization setting (refining a one-shot lead under a fixed compute budget) and compare MCTD-0 (random fill; no experts), MCTD-1 (single proposal expert), and MCTD-ME (multi-expert ensemble).
+Details on task-specific baselines/ablations are deferred to Appendix A.10.
+Datasets and splits.
+We evaluate on two benchmarks: CAMEO 2022 [12] and a PDB date-split [5].
+We follow the standard splits and filtering; statistics appear in Appendix A.2 (CAMEO: 183 targets, length $15/247.5/704$ min/avg/max; PDB date-split: a 449-target subset, length $2/242.2/511$).
+Ground-truth backbones and native sequences are available for both.
+Critics and evaluation metrics.
+All planner variants share the same task-specific composite critic; for each target we report Baseline, Final, and $\Delta$} (Final–Baseline) and aggregate by mean $\pm$ s.e.m., mapping scalars to $[0,1]$ (pLDDT residue-averaged and divided by $100$).
+During search, the critic uses only surrogate structure models/confidence scores (ground truth is used only for final reporting), and the critic is used only to rank/select edits.
+Full reward definitions, formulas, and weights appear in Appendix A.3, with further experimental setup details in Appendix A.10.
+5.2 Protein Folding
+We first evaluate MCTD-ME on the folding task: given an amino acid sequence, predict its 3D structure.
+We use the CAMEO 2022 benchmark [12] and a PDB date-split benchmark [5].
+As the initial lead, we take the one-shot DPLM-2 150M structure of the native sequence and measure improvement against the true structure (backbone RMSD, TM-score) in Table 1.
+We compare three variants: (i) Random (MCTD-0), which fills masked sites with random amino acids (no expert proposals); (ii) Single-Expert (MCTD-1), which uses exactly one pretrained DPLM-2 model (150M, 650M, or 3B) to generate rollouts; and (iii) MCTD-ME, which aggregates proposals from all three experts (DPLM-2 150M, 650M, 3B) inside a single tree.
+All variants share the same critic (ESMFold-based folding reward), pLDDT-guided masking schedule, and compute budget; only the proposal policy (number and type of experts and whether we use tree search vs.
+sampling) differs.
+On CAMEO we further include two ablations: Sampling (multi-expert) without tree search and MCTD-UCT (multi-expert) using standard UCB selection.
+Results are summarized in Table 1, which reports for each method the Baseline (one-shot ESMFold lead), the Final structure quality after planning, and the corresponding improvements ($\Delta$).
+Random search yields only small average gains, while incorporating learned DPLM-2 guidance substantially reduces RMSD and improves TM-score.
+On CAMEO, MCTD-ME attains large average gains over the baseline lead (RMSD drops from $10.50$ Å to $9.41$ Å, TM-score rises from $0.537$ to $0.827$), with final TM-score and composite reward at least matching, and often exceeding, the best single-expert configuration.
+On the PDB split, multi-expert MCTS again achieves strong lead refinement, with MCTS-based variants (especially $\text{P}\mathcal{H}\text{-UCT}$) substantially reducing RMSD from the baseline ($8.45$ Å to 6.476.476.736.73 Å) while raising TM-score from $0.591$ to $\sim 0.72$ and improving the composite reward.
+We also observe that a larger fraction of targets benefit from MCTD-ME compared to the best single expert, consistent with the per-configuration $\Delta$ columns in Table 1; CAMEO ablations further show that multi-expert Sampling improves RMSD but can hurt TM, whereas multi-expert MCTD-UCT improves both yet remains below MCTD-ME on TM and composite reward.
 5.3 Protein Inverse Folding
-We evaluate Monte Carlo Tree Diffusion with Multiple Experts (MCTD-ME) on the task of protein inverse folding, i.e.
-structure-conditioned sequence design.
-The main purpose of inverse folding is to find the amino acid sequence such that its folding matches a given backbone structure.
-In this work, we focus on this domain where ground truth sequences are known, leaving other applications (e.g.
-de novo structure generation) for future work.
-The evaluation uses the CAMEO 2022 dataset [10] and PDB data split ([4]), a standard benchmark of recently solved protein structures used in prior work.
-For each backbone structure in CAMEO, the goal is to generate an amino acid sequence that folds to that structure.
-We adopt the same metrics as DPLM-2 [31]: (1) Amino Acid Recovery (AAR) – the percentage of residues in the designed sequence that exactly match the native sequence (sequence recovery), and (2) self-consistency TM-score (scTM) – the TM-score measuring structural consistency between the target backbone and the structure predicted from the designed sequence.
-Higher AAR indicates better sequence accuracy for the given fold, while scTM reflects how well the designed sequence is likely to fold into the target structure.
-Following DPLM-2’s protocol, we report both metrics for the CAMEO test set.
-(For reference, co-generation baselines like MultiFlow [4] and ESM3 [6] attain  32–47% AAR on CAMEO, while the state-of-the-art diffusion model DPLM-2 (3B) achieves  52% AAR and 0.89 scTM)
-5.1.1 Baselines and Variants
-We conduct an ablation study to quantify the impact of multi-expert guidance. In particular, we compare three variants of our approach:
-MCTD-ME-0 (no experts): A control variant that performs the diffusion tree search with no expert guidance.
-The forward diffusion model’s proposals are accepted at random, yielding essentially a random-fill of the sequence (no reward-driven pruning).
-MCTD-ME-1 (single expert): Our MCTS-based diffusion planner using a single expert model as the critic.
-Here we use the DPLM-2 150M, 650M, 3B models seperately as the sole experts for scoring candidates (chosen for its strong performance and manageable runtime).
-This represents a single-criterion search.
-MCTD-ME-E (ensemble of experts): The full multi-expert MCTD using an ensemble of three DPLM-2 models of varying scales – 150M, 650M, and 3B parameters.
-These experts provide a diverse set of evaluations, which are combined into a composite reward signal.
-All variants use the same MCTS sampling pipeline described in Section 4.
-At each diffusion step (node expansion), the planner simulates rollouts to incrementally denoise the sequence.
-In our implementation, each expert performs 2 rollouts, and with 3 experts this yields around 6 candidate expansions per node.
-To limit branching, we retain the Top-2 candidates according to the multi-critic reward at each step.
-The MCTS selection policy uses a UCT-style formula adapted for multi-expert value estimates with an exploration constant $c_{p}=1.414$.
-We also incorporate an entropy-aware bonus as described (Section 4.3), so that nodes with high disagreement among experts are explored more.
-All expert models are pre-trained (no further fine-tuning); MCTD operates as a test-time algorithm that plugs these experts into the search.
-Reward Formulation: The multi-objective reward $R$ used to guide MCTD-E is a weighted sum of the expert criteria.
-Specifically, we define $R=0.6\text{AAR}+0.35\text{scTM}+0.05B$, where B is a small bonus from simple biophysical properties (average hydrophobicity, amino acid composition frequency, and net charge).
-The 60/35/5 weighting was chosen to prioritize sequence recovery while still strongly rewarding structural fidelity, plus a minor regularization to favor realistic amino acid distributions.
-This composite reward design is in line with prior multi-critic strategies.
-In our setting, AAR and scTM are computed after a full sequence is generated (i.e.
-at a leaf node of the search); partial sequences during tree expansion receive interim estimates (for example, we estimate scTM for a partial sequence by filling missing residues with the most likely amino acids before structure prediction, and likewise compute a partial AAR by comparing the filled-in sequence to the native).
-These approximations guide the search towards promising candidates even before the sequence is complete.
-5.1.2 Results
-Table 1 summarizes the inverse folding results on the CAMEO dataset.
-Overall, MCTD-ME-E (multi-expert) achieves the best performance, outperforming both the single-expert and no-expert variants by a significant margin.
-MCTD-ME-E attains an average AAR of 44.5%, higher than the 43.2% achieved by the strongest single-expert run (650M, expert 0).
-The normalized reward of MCTD-ME-E designs averages 0.303, also exceeding the 0.295 of the 650M expert alone.
-By contrast, the MCTD-ME-1 (150M expert, expert 1) variant essentially matches its baseline (42.7% AAR, 0.290 reward), while the MCTD-ME-1 (3B expert, expert 2) variant actually degrades performance (37.0% AAR, 0.254 reward).
-This indicates that our MCTS planner by itself introduces little overhead or degradation when only one expert is used – essentially reproducing the single model’s capabilities – but the ensemble of experts (MCTD-ME-E) yields a further boost: +1.3 percentage points in sequence recovery and a notable gain in normalized reward relative to the best single expert.
-This suggests that the ensemble is successfully steering the generation to find sequences that satisfy multiple criteria better than any single model alone.
-As expected, MCTD-ME-0 (random fill) performs near chance-level – we observed 38.7% AAR (roughly the random identity baseline for 20 amino acids) and very low reward (0.272), confirming that unguided diffusion yields essentially arbitrary sequences with poor folding to the target.
-These serve as a sanity check that improvements of MCTD-ME-1 and MCTD-ME-E are indeed due to learned guidance rather than artifact.
-Importantly, MCTD-ME-E not only improves average performance but also reduces failure cases.
-We found that MCTD-ME-E produced valid, foldable sequences for nearly all test structures, whereas single-expert runs occasionally returned sequences with poor recovery or structural inconsistencies.
-The multi-expert approach mitigates this by combining perspectives: the smaller 150M expert, for instance, tends to propose more diverse amino acid choices (higher entropy), while the larger 3B expert strongly evaluates overall fitness.
-The MCTS search benefits from this diversity–exploitation mix, as reflected in the higher overall reward of MCTD-ME-E designs.
-We also observed that the advantage of multiple experts grows with problem complexity.
-On longer proteins (length >300), MCTD-ME-E’s sequence recovery was several percentage points higher than MCTD-ME-1 on average, whereas on short proteins (<100 residues) both methods performed similarly.
-This trend aligns with the intuition that larger search spaces and more complex structures gain more from an ensemble of critics guiding the search.
-In sum, multi-expert MCTD consistently outperforms both single-expert planning and one-shot diffusion decoding on this task, achieving state-of-the-art results on the CAMEO inverse folding benchmark (to our knowledge).
-See Fig 3 for length-binned improvements in AAR, reward, and scTM.
-Figure 3: Comparison of three evaluation metrics (AAR, normalized reward, and scTM) across different model variants.
-Higher values indicate better sequence recovery and structural consistency, showing the advantage of multi-expert MCTD.
+Variant Motif RMSD $\downarrow$ (Å) scTM $\uparrow$ Reward $\uparrow$ Success $\uparrow$
+Base Final $\Delta$ Base Final $\Delta$ Base Final $\Delta$ Rate
+Baseline (DPLM-2 650M) $\mathbf{2.971\pm 3.278}$ $\mathbf{0.717\pm 0.129}$ $\mathbf{0.433\pm 0.185}$ 10/23
+DPLM2-650M (Single-Expert) $2.388\pm 2.943$ $0.583\pm 0.545$ $747\pm 0.148 +0.0$$030\pm 0.022$ $460\pm 0.182 +0.0$$026\pm 0.016$ 14/23
+RFDiffusion (Single-Expert) $1.501\pm 1.659$ $1.470\pm 0.876$ $631\pm 0.150 0.0$$086\pm 0.027$ $446\pm 0.181 +0.0$$013\pm 0.019$ 16/23
+FoldFlow (Single-Expert) $3.874\pm 2.278$ $0.903\pm 0.800$ $641\pm 0.144 0.0$$076\pm 0.028$ $408\pm 0.129 0.0$$025\pm 0.013$ 8/23
+ProteinA (Single-Expert) $1.313\pm 1.148$ $1.658\pm 1.011$ $638\pm 0.154 0.0$$079\pm 0.020$ $450\pm 0.140 +0.0$$017\pm 0.010$ 14/23
+MCTD-UCT (Multi-Expert) $3.430\pm 1.30$ $0.46\pm 0.39$ $730\pm 0.053 +0.0$$014\pm 0.018$ $438\pm 0.043 +0.0$$005\pm 0.011$ 13/23
+Sampling (Multi-Expert) $2.581\pm 1.17$ $0.39\pm 0.90$ $\mathbf{0.759\pm 0.136}$ $\mathbf{+0.042\pm 0.027}$ $482\pm 0.064 +0.0$$049\pm 0.026$ 15/23
+MCTD-ME (Multi-Expert) $\mathbf{1.493\pm 1.675}$ $\mathbf{+1.478\pm 1.453}$ $742\pm 0.049 +0.0$$025\pm 0.120$ $\mathbf{0.510\pm 0.187}$ $\mathbf{+0.078\pm 0.042}$ 17/23
+Table 3: Motif scaffolding results (mean $\pm$ s.e.m.; n=23n{=}23 motifs).
+The first row reports the Baseline from one-shot DPLM-2 (650M).
+Variant rows show the identical baseline values (left columns), the Final performance after MCTD-ME refinement, and the improvement $\Delta$ (for RMSD, $\Delta{=}\text{Baseline}{-}\text{Final}$ in Å; for scTM/Reward, $\Delta{=}\text{Final}{-}\text{Baseline}$).
+We compare (i) single-expert planners (four rows), (ii) a multi-expert Sampling baseline (no tree search; experts directly propose on pLDDT masks), (iii) multi-expert UCT (MCTS with standard UCB selection), and (iv) our full MCTD-ME (multi-expert with uncertainty/novelty-aware selection).
+Success is defined as Final motif-RMSD <1<1Å and scTM >0.8>0.8, and the Success Rate is the fraction of motifs meeting this criterion.
+The DPLM-2 baseline here is one-shot generation; additional sampling-based results are reported in Appendix A.8.
+We evaluate MCTD-ME on structure-conditioned sequence design (CAMEO 2022 and a PDB date-split with known natives), reporting amino-acid recovery (AAR; fraction matching native) and self-consistency TM-score (scTM; TM between the target backbone and the fold of the designed sequence) following Wang et al.
+[37].
+We ablate proposal experts: MCTD-0 (no experts; random fills), MCTD-1 (single proposal expert; separate runs with DPLM-2 150M, DPLM-2 650M, and ProteinMPNN [7]), and MCTD-ME (ensemble of those experts).
+In addition, we include two multi-expert ablations on CAMEO only: Sampling (Multi-Expert) without tree search, and MCTD-UCT (Multi-Expert) using standard UCB selection.
+All proposals are scored by the same critic with reward $R=0.60\,\text{AAR}+0.35\,\text{scTM}+0.05\,B$ (small biophysical bonus $B$), and sequences are completed via masked-diffusion rollouts within MCTS.
+As shown in Table 2, MCTD-ME consistently improves structural fidelity and overall reward: on both CAMEO and PDB it delivers the largest scTM and normalized-reward gains over baselines, while AAR is maintained on CAMEO and increases on PDB.
+Random edits (MCTD-0) degrade AAR and yield only small structural gains, confirming the need for guided proposals.
+Single-expert planners track their proposal models—e.g., 150M preserves its baseline; 650M improves both scTM and AAR—while ProteinMPNN alone can trade off AAR for scTM.
+The multi-expert ablations show partial benefits: Sampling captures some gains without planning, and UCT recovers additional benefit with tree search, but both remain below MCTD-ME, underscoring the value of uncertainty/novelty-aware selection.
+The ensemble’s advantage grows on harder cases (e.g., longer proteins), reflecting benefits from expert diversity under a common critic.
+See Appendix A.4 (Fig.
+3) for AAR, reward, and scTM improvements across protein-length bins.
+5.4 Motif Scaffolding
+Finally, we test conditional generation-motif scaffolding: given a functional motif (its amino acid sequence and 3D structure), generate the remaining scaffold sequence and structure around it.
+Following prior work (e.g., FrameFlow [41]), we use a PDB-sourced benchmark (EvoDiff’s $\sim$24 curated motifs[42]).
+The baseline model is DPLM-2 (150M) run in conditional mode to co-generate scaffold sequence/structure given the motif.
+Our planner uses the same MCTS budget and masking schedule as in other tasks (Appendix A.10): at each expansion, only scaffold positions are masked and proposed.
+We ablate proposal policies and expert sets: single-expert planners (DPLM-2 650M, RFDiffusion [39], ProteinA [9], FoldFlow [4]), Multi-Expert Sampling (no tree search; experts directly propose on pLDDT masks), Multi-Expert UCT (standard UCB selection), and MCTD-ME (multi-expert with uncertainty/novelty-aware selection).
+Across all variants, candidates are scored by the same composite critic, and the motif residues are hard-frozen throughout generation.
+Metrics are: motif-RMSD (RMSD of the designed scaffold after rigidly aligning on the motif), scTM (TM-score between the designed full backbone and native), and mean pLDDT over scaffold residues.
+Table 3 summarizes results on the EvoDiff curated motif set (n=23n{=}23; we exclude 4JHW because its PDB is missing chain A, preventing motif extraction with the standard preprocessing).
+Overall, MCTD-ME achieves the strongest aggregate performance, reducing motif-RMSD from $2.971$ to $1.493$ Å while improving reward from $0.393$ to $0.481$ and increasing success from $10/23$ to $17/23$.
+Compared to single-expert planners, MCTD-ME is the only setting that simultaneously maintains low motif-RMSD (comparable to the best structure-only expert) while avoiding the scTM degradation seen in some specialists.
+The Sampling (Multi-Expert) ablation recovers a substantial fraction of the reward/scTM gains (it attains the best scTM), but yields weaker motif fidelity and a lower success rate, consistent with the absence of sequential credit assignment.
+The MCTD-UCT ablation recovers some of the planning benefit over Sampling but remains below MCTD-ME in both motif-RMSD and overall reward, highlighting the importance of uncertainty/novelty-aware expert routing.
+Overall, aggregating diverse experts under a shared critic with informed selection mitigates the per-expert trade-offs, yielding robust improvements across motifs while strictly preserving the motif.
 6 Conclusion
-In conclusion, we introduced MCTD-ME, which combines diffusion-based generative planning with MCTS and multi-expert guidance.
-The method is novel in reframing diffusion decoding as tree search for parallel multi-token exploration and in integrating multiple expert evaluators with a pLDDT-guided masking schedule to improve biophysical fidelity.
-Our selection policy ($\text{P}\mathcal{H}\text{-UCT}\text{-ME}$) efficiently arbitrates among experts using cached uncertainty statistics, enabling stable and low-overhead rollouts.
-Empirically, MCTD-ME outperforms state-of-the-art baselines on inverse folding, scales effectively with additional computation, and provides a general template for diffusion-plus-search workflows in structured generative tasks.
+We presented MCTD-ME, a model-agnostic planner that reframes reverse diffusion as Monte Carlo tree search, enabling parallel multi-token revisions and principled exploration guided by multiple proposal experts.
+A pLDDT-driven masking schedule focuses denoising on structurally uncertain regions, improving biophysical fidelity, while our $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$ selection reuses cached uncertainty to efficiently arbitrate among experts without recomputing entropies.
+Across folding, inverse folding, and motif scaffolding, MCTD-ME consistently outperforms single-expert and one-shot baselines—raising AAR and scTM and lowering RMSD—with gains that scale with test-time compute and are robust to the choice of experts.
+Although the method currently relies on surrogate structure predictors and incurs some search overhead as a limitation, it provides a unified, plug-and-play test-time mechanism that reliably improves design quality.
+Future work will integrate learned critics and apply further systems optimizations to enhance both accuracy and speed.
 Appendix A Appendix
-A.1 Algorithm details
+A.1 Algorithm details.
+Lead optimiaztion
+In this algorithm, we use MCTD-ME to tackle the lead optimization challenge.
+Beginning with an initial lead—a fully denoised baseline sequence—we execute MCTS with the following procedure: (i) selection uses a $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$–style rule that combines value (Q), exploration, and cached uncertainty bonuses (ensemble surprisal ($\mathcal{U}_{\text{ent}}$) and novelty ($\mathcal{U}_{\text{div}}$)); (ii) expansion proposes edits only on pLDDT-masked uncertain sites and performs multi-expert rollouts (E generators), producing candidates scored once via a composite critic (cached to avoid re-evaluation); (iii) we keep the Top-K children by composite reward, attach their uncertainty bonuses, and (iv) backpropagate the candidate’s value to update ancestors (max/sum backup).
+This yields a planner that routes among experts, focuses edits where structure is uncertain, reuses evaluations via caching, and optimizes a task-aware composite objective.
 Algorithm 1 Monte Carlo Tree Diffusion with Multiple Experts (MCTD-ME)
-1:
-[…]
-[…]
-4:$K$: number of children per expansion
-5:$E$: number of expert models
-6:$R$: rollouts per expert
-7:$T$: total simulations
 […]
@@
 […]
-11:for $i=1$ to $T$ do
-12:  $node\leftarrow root$
-13:  $\triangleright$ /* Selection via $\text{P}\mathcal{H}\text{-UCT}\text{-ME}*$/
-14:  while $node.children\neq\emptyset$ do
-15:   $node\leftarrow\text{P}\mathcal{H}\text{-UCT}\text{-ME}(node.children)$ // uses $Q$, UCB, and cached $(\mathcal{U}_{\text{ent}},\mathcal{U}_{\text{div}})$
-16:  $\triangleright$ /* Expansion via pLDDT Masking and Multi-Expert Rollouts */
-17:  $\mathcal{M}\leftarrow\textsc{GetMaskSet}(node.sequence)$ // Based on pLDDT
-18:  $\mathcal{Y}\leftarrow\emptyset$ // candidate children
-19:  for $e\in\mathcal{E}$ do // each generator expert
-20:   for $r=1$ to $R$ do // rollouts
-21:     $y_{0}\leftarrow\textsc{MaskedDiffusion}(node.sequence,\mathcal{M},e)$
-22:     if $y_{0}\notin cache$ then
-23:      $score\leftarrow\textsc{EvalComposite}(y_{0};\mathcal{C})$ // weighted sum of critics
-24:      $cache[y_{0}]\leftarrow score$
-25:     $\mathcal{Y}\leftarrow\mathcal{Y}\cup\{y_{0}\}$
-26:  $\mathcal{Y}_{\text{top}}\leftarrow\textsc{TopK}(\mathcal{Y},K;\,cache)$ // sort by composite reward
-27:  for $y^{\prime}\in\mathcal{Y}_{\text{top}}$ do
-28:   $(\mathcal{U}_{\text{ent}}(y^{\prime}),\mathcal{U}_{\text{div}}(y^{\prime}))\leftarrow\textsc{ComputeBonuses}(y^{\prime},\,node.sequence,\,\mathcal{M})$ // ensemble surprisal & novelty
-29:   $\textsc{AddChild}(node,y^{\prime},\,\mathcal{U}_{\text{ent}}(y^{\prime}),\,\mathcal{U}_{\text{div}}(y^{\prime}))$ // create new node with full seq
-30:  $\triangleright$ /* Evaluation of New Children (full-sequence) */
-31:  for $y^{\prime}\in\mathcal{Y}_{\text{top}}$ do
-32:   $v\leftarrow cache[y^{\prime}]$ // may hit cache; same composite reward as above
-33:   $\triangleright$ /* Backpropagation */
-34:   $\textsc{Backpropagate}(y^{\prime},\,v)$ // supports max or sum backup
-35:return Top-$k$ sequences by $cache[\cdot]$
-A.2 Dataset Details
+[…]
+De novo generation
+We start from an all-mask root.
+At each iteration, Selection ($\text{P}\mathcal{H}\text{-UCT}\text{-ME}$) walks to a leaf.
+Expansion proposes partial fills only at low-pLDDT sites (high-pLDDT positions are frozen).
+Then Evaluation performs short rollouts from each new child (e.g., a few masked-diffusion steps / fast fold+critic passes) to estimate a terminal composite value; results are cached.
+Finally, we Backpropagate those values.
+This keeps the planner focused on uncertain regions while using cheap rollouts to decide which partials to keep; see Algorithm 2 for pseudocode.
+Algorithm 2 De Novo MCTD with Multiple Experts (MCTD-ME -DeNovo)
+[…]
+[…]
+[…]
+A.2 Dataset details.
 We evaluate MCTD-ME on two protein sequence benchmarks at the chain level.
-(1) CAMEO2022 comprises 183 targets released in 2022; we use the reference backbones and evaluate inverse folding on the corresponding chains.
+As illustrated in Table 4 (1) CAMEO2022 comprises 183 targets released in 2022; we use the reference backbones and evaluate inverse folding on the corresponding chains.
 (2) PDB_date is a date-split benchmark of 7,855 proteins constructed from the Protein Data Bank (PDB); we hold out targets by deposition date and evaluate inverse folding on their chains.
@@
 Note: Experiments use a date-split subset of 449 PDB_date chains; its length distribution is similar to the full set.
-A.3 Case Study on CAMEO data
-Fig.
-4 overlays three structures to illustrate how MCTD-ME steers sequences toward the target fold.
-The red ribbon is the CAMEO reference (ground truth), the blue ribbon is the structure predicted from the baseline sequence, and the multicolor ribbon is our MCTD-ME design.
-For the multicolor model, residues are shaded on a red $\rightarrow$ blue gradient proportional to per-residue deviation from the reference (red = near, blue = far), making local improvements visually apparent along the chain.
-Relative to the blue baseline, the MCTD-ME design shows widespread “reddening,” especially across helices and connecting loops, indicating closer alignment to the target geometry while preserving already well-formed secondary structure.
-This behavior matches the algorithm’s intent: pLDDT-aware masking focuses edits on uncertain segments, while multi-expert rollouts—selected via $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$—propose diverse fixes that are promoted during search, yielding higher structural fidelity (e.g., scTM) without over-editing stable regions.
-Figure 4: CAMEO 7dz2_C overlay.
-Red: target reference.
-Blue: baseline (starting point).
-Multicolor: MCTD design, colored per residue from red (near reference) to blue (far), after structural alignment.
-The shift from blue toward red across many regions visualizes how MCTD moves the baseline toward the reference via pLDDT-aware masking and multi-expert guided rollouts.
-A.4 Setup and Hyperparameters
-For equitable assessment, we compare all variants under matched search and decoding budgets.
-Unless otherwise specified, every method uses the same DPLM-2 decoding configuration (deterministic unmasking with argmax sampling) and identical reward evaluation.
-In each table/figure, competing methods share the same rollout budget and selection rule; only the availability of diffusion experts differs across modes (random_no_expert, single_expert, multi_expert).
-We conducted only light tuning around a small grid and then fixed a single setting for all reported comparisons.
+A.3 Reward definitions and normalization.
+Terminology (metrics).
+TM-score: standard topology-similarity score between two backbones (range $[0,1]$; higher is better).
+scTM: “sequence-conditioned TM-score” — TM-score between the target backbone and the backbone predicted from the designed sequence (higher is better).
+[…]
+[…]
+pLDDT: per-residue confidence from a structure predictor (e.g., AlphaFold-type models), originally reported on a $[0,100]$ scale; we divide by $100$ to normalize to $[0,1]$.
+AAR: amino-acid recovery — fraction of positions where the designed sequence exactly matches the native (range $[0,1]$).
+Composite reward: weighted combination of normalized terms (e.g., TM-score, transformed RMSD, pLDDT, scTM, biophysical checks), constructed to lie in $[0,1]$.
+All reward components are normalized to the $[0,1]$ range; pLDDT is averaged over residues and divided by $100$. We use the composite reward score as follows:
+[…]
+$R_{\mathrm{fold}}=\alpha\,\mathrm{TM}+\beta\,\bigl(1-\min(\mathrm{RMSD}/10,\,1)\bigr)+\gamma\,\overline{\mathrm{pLDDT}},$
+[…]
+[…]
+$R_{\mathrm{inv}}=0.60\,\mathrm{AAR}+0.35\,\mathrm{scTM}+0.05\,B,$
+where AAR is residue-wise recovery versus the native sequence, scTM is the TM-score between the target backbone and the fold predicted from the designed sequence, and $B$ is a small biophysical bonus (e.g., motif/chemistry checks).
+[…]
+Candidates that do not exactly preserve the motif receive $R_{\mathrm{motif}}{=}0$. Otherwise,
+$R_{\mathrm{motif}}=0.40\,\overline{\mathrm{pLDDT}}_{\mathrm{scaf}}+0.30\,g(\mathrm{RMSD}_{\mathrm{motif}})+0.30\,\mathrm{scTM}+0.20\,\mathbf{1}\!\big[\mathrm{RMSD}_{\mathrm{motif}}{<}1\wedge\mathrm{scTM}{>}0.8\big],$
+[…]
+$g(x)=\begin{cases}\max(0,\,1-x/2),&x<1,\\[2.0pt] \max(0,\,0.2-x/10),&x\geq 1.\end{cases}$
+We use a broad inverted normalization for global folding RMSD and a sharper piecewise map (plus a success bonus) for local motif RMSD to prioritize precise interface preservation.
+A.4 Length-binned improvements.
+Here we stratify inverse-folding gains by protein length to assess scalability.
+Figure 3 reports mean improvements (Final–Baseline) for AAR, normalized reward, and scTM across five bins: <<100, 100–200, 200–300, 300–400, and >>400 residues.
+The multi-expert planner outperforms the single-expert baseline in every bin, with especially pronounced margins on long proteins: scTM gains remain positive and grow with length (single-expert even turns slightly negative in the 300–400 bin), reward boosts are largest for >>400, and AAR gains—while modest—are consistently higher for the ensemble (peak around 200–300).
+These trends indicate that expert diversity is most beneficial as sequence–structure complexity increases.
+AAR is a strict token-identity metric and thus less sensitive to small but structurally meaningful edits, whereas scTM/reward are smoother structural objectives that respond strongly to localized corrections.
+In addition, our critic emphasizes structural fidelity and the pLDDT-guided masking focuses changes on low-confidence regions, encouraging fold-improving (not necessarily native-identical) substitutions—producing larger scTM/reward deltas than AAR, especially for longer, more complex proteins.
+[…]
+A.5 Lead-optimization vs. De novo generation.
+While our current implementation treats root node as a complete sequence—effectively a lead-optimization style search—the framework naturally generalizes to fully de novo generation.
+In this setting, the root node could be a fully masked sequence, with subsequent nodes representing partially unmasked intermediates obtained through pLDDT-guided masking.
+Evaluation would then involve a simulation stage for reward estimation at each partial completion.
+Conceptually, this is equivalent to our guided unmasking scheme, differing only in granularity: entropy and disagreement are currently assessed at the level of entire candidate sequences, but under de novo generation they would be localized to the partially unmasked subsequence.
+As a result, performance is expected to be qualitatively similar, with the search dynamics adapting seamlessly from complete-sequence optimization to progressive de novo exploration.
+A.6 Related works.
+Protein and Molecular Diffusion Models.
+Recent advances apply diffusion models to protein design across both structure and sequence domains.
+RFdiffusion [38] introduced a guided 3D diffusion model for protein backbones.
+On the sequence side, DPLM [36] and its extension DPLM-2 [37] demonstrated that discrete diffusion pretraining on large-scale protein data enables high-quality sequence generation and versatile conditioning (e.g., inverse folding, property steering).
+Most recently, DiffGui [10] proposed a target-aware, E(3)-equivariant diffusion framework for joint atom and bond generation, incorporating multi-property guidance (e.g., binding affinity, drug-likeness) to generate realistic, high-affinity 3D molecules.
+However, these diffusion approaches generally rely on one-shot or open-loop decoding, making it difficult to incorporate adaptive search or revision during generation—highlighting the need for planning methods such as MCTS.
+Monte Carlo Tree Search in Generative Modeling.
+Monte Carlo Tree Search (MCTS) has improved generative decoding in tasks requiring complex reasoning or program synthesis.
+In code generation, AlphaCode [16] used a brute-force generate-and-test method by sampling large program sets and selecting those that passed tests.
+More structured methods followed: RethinkMCTS [15] performs MCTS over LLM reasoning steps, using execution feedback to refine intermediate thoughts.
+Tree-of-Thoughts [40] frames reasoning as a tree of self-evaluated partial solutions expanded via classical search (DFS/BFS), boosting performance on puzzles and planning.
+PG-TD [43] simulates lookahead in code generation by executing candidate programs during decoding.
+ERP [23] achieves state-of-the-art performance by leveraging MCTS with a GPT model, with applications in both drug discovery and code generation.
+Across domains, MCTS-style planning consistently enhances generation quality and reliability.
+Yet, most of these approaches are tied to autoregressive GPT-style models, which remain token-by-token and struggle with long-range dependencies—suggesting that discrete diffusion could provide a more flexible backbone for planning.
+Discrete Diffusion for Planning and Generation.
+While originally developed for continuous domains, diffusion models have been successfully adapted to discrete sequence generation and planning.
+D3PM [1] introduced absorbing-state (masking) noise for discrete diffusion, achieving competitive text generation.
+Follow-ups diffusion language models like MD4 [32] further closed the performance gap with autoregressive models.
+In planning, Diffuser [11] modeled offline RL as trajectory denoising, generating full state-action paths guided by value functions.
+To enable adaptive inference, Monte Carlo Tree Diffusion (MCTD) [2] reimagines diffusion as tree search, branching multiple denoising outcomes per step and scoring partial trajectories via heuristics.
+In protein design, ProtInvTree [20] similarly integrates reward-guided search with jumpy denoising for efficient inverse folding.
+More recently, PepTune [34] proposes multi-objective discrete diffusion for therapeutic peptide SMILES, pairing masked diffusion with an MCTS-style guidance strategy to negotiate tradeoffs among binding, solubility, and membrane permeability objectives.
+Still, existing methods optimize toward a single diffusion generator, which limits their ability to effective search space exploration in biological and chemical design.
+Multi-Experts Guided Generation.
+Multi-expert systems leverage several generative models, or ’experts,’ in tandem, allowing each to contribute its strengths while combining their state-wise expertise.
+This idea has been explored in reinforcement learning [21, 22] and bandit [24] settings, where algorithms learn from multiple expert policies or oracles to improve decision-making.
+In generative modeling, mixture-of-experts strategies similarly combine specialized models to better cover complex data distributions.
+For example, recent diffusion-based image generators use multiple specialized denoising models across different noise levels [8] or even heterogeneous model architectures [14] to boost output quality.
+However, these multi-expert approaches remain largely static and one-shot during inference, without an adaptive search process.
+To our knowledge, no prior diffusion model integrates multiple experts into a planning loop for generative design.
+By incorporating a team of expert generators within an MCTS-driven diffusion framework, our work enables dynamic collaboration among experts to satisfy multiple design criteria and explore a broader solution space—an innovative combination that extends diffusion-based generation to more complex, multi-objective scenarios.
+Complementary to search-based methods, DRaFT and ReFT-style approaches optimize generation via training-time reward alignment or parameter-efficient finetuning [6, 28]; in contrast, our framework is training-free and operates at test time by planning over diffusion rollouts.
+A.7 Case study on CAMEO data and motif scaffolding.
+Figure 4 shows how MCTD-ME refines a CAMEO target (7dz2_C).
+The light-gray ribbon is the baseline lead (one-shot structure).
+Over it, the MCTD-ME design is drawn only where an improvement over the lead is made and is colored by per-residue closeness to the native backbone after alignment: red = very close, yellow = closer but not as close.
+(Regions without improvement are left in gray and reflect the baseline.) We observe broad reddening across helices and connecting loops, indicating tighter agreement with the target while preserving segments the lead already modeled well.
+This aligns with our mechanism: pLDDT-aware masking concentrates edits on low-confidence sites, and PH-UCT-ME promotes diverse, high-value proposals from multiple experts, yielding higher fold consistency without over-editing stable regions.
+Figure 4: CAMEO 7dz2_C refinement.
+Light gray: baseline lead.
+Colored overlay: MCTD-ME edits where improvement is achieved, shaded by closeness to the native backbone (red = very close; yellow = closer but not as close).
+Gray segments indicate no change from the lead.
+The predominance of red/yellow across many regions visualizes how MCTD-ME moves the lead toward the native geometry via pLDDT-aware masking and multi-expert selection.
+Figure 5 visualizes how MCTD-ME refines the scaffold around a fixed functional motif.
+The light, semi-transparent segment indicates the input motif (hard-frozen throughout generation).
+The surrounding gray ribbon is the baseline scaffold (lead), while the colored overlay shows MCTD-ME edits only where improvement over the lead is achieved, shaded by closeness to the native structure after motif-aligned superposition (red = very close, yellow = closer but not as close).
+We observe concentrated reddening along the $\beta$-strand and adjacent loop, indicating better register and loop geometry at the motif–scaffold interface, while unchanged regions remain gray—highlighting that MCTD-ME preserves already adequate geometry and focuses edits where they matter.
+Figure 5: Motif scaffolding (PDB 7mrx).
+Light/white segment: fixed input motif; gray ribbon: baseline scaffold (lead).
+Colored overlay: MCTD-ME ’s improved scaffold residues (red = very close to native; yellow = closer), shown only where the scaffold improves relative to the lead after motif-aligned superposition.
+The localized reddening near the motif indicates targeted refinement of the interface without over-editing stable regions.
+A.8 Additional experiments.
+A.8.1 Random-lead evaluation (model-agnosticity).
+To complement the fixed-lead analysis, we report inverse-folding results when the baseline is a fresh one-shot sample per target (“random lead”) from DPLM-2 (150M), introducing stochastic variation across runs (see Table 5).
+Despite this added baseline noise, MCTD-ME consistently attains the strongest improvements in normalized reward and scTM, while maintaining (CAMEO) or improving (PDB) AAR relative to single-expert controls and random routing.
+Taken together with the fixed-lead results in the main text, these outcomes demonstrate model agnosticity: the planner reliably refines any provided lead—whether fixed or randomly resampled—under a shared critic.
+Minor differences in $\Delta$ magnitudes reflect baseline variability rather than changes to the optimization policy.
+Variant AAR $\uparrow$ Normalized Reward $\uparrow$ scTM $\uparrow$
+Baseline Final $\Delta$ Baseline Final $\Delta$ Baseline Final $\Delta$
+CAMEO Benchmark
+MCTD-0 (Random) $0.446\pm 0.011$ $387\pm 0.009 0.0$$058\pm 0.004$ $0.417\pm 0.009$ $416\pm 0.008 0.0$$001\pm 0.003$ $0.378\pm 0.013$ $440\pm 0.014 +0.0$$062\pm 0.009$
+Single-Expert (150M) $0.441\pm 0.009$ $441\pm 0.009 +0.0$$000\pm 0.000$ $0.289\pm 0.008$ $289\pm 0.008 +0.0$$000\pm 0.000$ $0.371\pm 0.015$ $371\pm 0.015 +0.0$$000\pm 0.000$
+Single-Expert (650M) $0.443\pm 0.009$ $458\pm 0.009 +0.0$$015\pm 0.002$ $0.435\pm 0.011$ $430\pm 0.010 0.0$$005\pm 0.004$ $0.371\pm 0.015$ $443\pm 0.015 +0.0$$072\pm 0.008$
+Single-Expert (ProteinMPNN) $0.448\pm 0.013$ $392\pm 0.012 0.0$$056\pm 0.005$ $0.430\pm 0.011$ $422\pm 0.010 0.0$$007\pm 0.004$ $0.366\pm 0.016$ $448\pm 0.015 +0.0$$082\pm 0.009$
+MCTD-ME (Multi-Expert) $\mathbf{0.452\pm 0.013}$ $\mathbf{0.445\pm 0.012}$ $\mathbf{-0.007\pm 0.005}$ $\mathbf{0.432\pm 0.012}$ $\mathbf{0.474\pm 0.011}$ $\mathbf{+0.041\pm 0.006}$ $\mathbf{0.342\pm 0.018}$ $\mathbf{0.445\pm 0.018}$ $\mathbf{+0.103\pm 0.013}$
+PDB Benchmark
+MCTD-0 (Random) $0.434\pm 0.003$ $401\pm 0.003 0.0$$033\pm 0.002$ $0.213\pm 0.002$ $200\pm 0.002 0.0$$013\pm 0.001$ $0.355\pm 0.003$ $376\pm 0.003 +0.0$$021\pm 0.002$
+Single-Expert (150M) $0.460\pm 0.006$ $460\pm 0.006 +0.0$$000\pm 0.000$ $0.290\pm 0.006$ $290\pm 0.006 +0.0$$000\pm 0.000$ $0.370\pm 0.007$ $370\pm 0.007 +0.0$$000\pm 0.000$
+Single-Expert (650M) $0.508\pm 0.007$ $541\pm 0.007 +0.0$$033\pm 0.003$ $0.348\pm 0.007$ $382\pm 0.006 +0.0$$034\pm 0.003$ $0.392\pm 0.008$ $426\pm 0.008 +0.0$$034\pm 0.003$
+Single-Expert (ProteinMPNN) $0.451\pm 0.016$ $398\pm 0.016 0.0$$053\pm 0.006$ $0.330\pm 0.013$ $318\pm 0.012 0.0$$012\pm 0.004$ $0.384\pm 0.009$ $444\pm 0.009 +0.0$$060\pm 0.005$
+MCTD-ME (Multi-Expert) $\mathbf{0.504\pm 0.006}$ $\mathbf{0.578\pm 0.006}$ $\mathbf{+0.074\pm 0.004}$ $\mathbf{0.353\pm 0.007}$ $\mathbf{0.427\pm 0.007}$ $\mathbf{+0.074\pm 0.004}$ $\mathbf{0.383\pm 0.007}$ $\mathbf{0.448\pm 0.008}$ $\mathbf{+0.065\pm 0.004}$
+Table 5: Inverse folding results on CAMEO and PDB.
+Baseline sequences are generated by one-shot DPLM-2 (150M) , which serves as the initialization for our lead-optimization procedure.
+Baselines may vary slightly across runs (see baseline-invariant results in Appendix A.8), while Final denotes the sequence after MCTD-ME refinement.
+Metrics reported include amino acid recovery (AAR), normalized reward, and scTM, with $\Delta$ indicating improvements over baseline.
+A.8.2 One-shot expert baselines (no refinement)
+Inverse folding (one-shot experts).
+Table 6 reports one-shot inverse-folding quality for each expert generator without any refinement.
+DPLM-2 scales favorably from 150M to 650M across AAR, scTM, and the composite metric; ProteinMPNN trades higher AAR for stronger biophysical plausibility (reflected in the composite).
+For a fair comparison to these single-shot experts, we also include a de novo variant of MCTD-ME (MCTD-ME) that samples a single sequence per target using the same stochastic rollout settings.
+Note this setting differs from our lead-optimization regime used elsewhere; results here should therefore not be conflated with tables that evaluate refinement starting from a fixed lead.
+Method AAR (mean $\pm$ std) [min–max] scTM (mean $\pm$ std) [min–max] Composite (mean $\pm$ std) [min–max]
+DPLM-2 150M (stoch) $370\pm 0.114 [0.0$31–0.799] $321\pm 0.061 [0.0$87–0.490] $0.381\pm 0.080$ [0.136–0.641]
+DPLM-2 650M (stoch) $403\pm 0.128 [0.0$51–0.824] $326\pm 0.060 [0.0$93–0.501] $0.402\pm 0.090$ [0.167–0.658]
+ProteinMPNN (stoch) $242\pm 0.064 [0.0$29–0.360] $0.309\pm 0.063$ [0.118–0.498] $0.300\pm 0.054$ [0.107–0.387]
+MCTD-ME $432\pm 0.119 [0.0$82–0.827] $0.329\pm 0.056$ [0.124–0.497] $0.420\pm 0.081$ [0.200–0.659]
+Table 6: Inverse folding (CAMEO, n=183n{=}183) with stochastic one-shot baselines plus MCTD-ME de novo.
+Each expert baseline emits a single sequence per target without refinement, using stochastic sampling configured to mirror MCTS rollouts: --temperature=1.0, --unmasking_strategy=stochastic1.0, --sampling_strategy=annealing@2.2:1.0.
+We report amino-acid recovery (AAR), self-consistency TM-score (scTM), and the composite reward as mean $\pm$ std across targets, with min–max in brackets.
+MCTD-ME (MCTS) is shown for reference (refined result under planning).
+Forward folding (one-shot structures).
+As a complementary check of generative fidelity, Table 7 evaluates one-shot forward folding (single prediction per sequence; no refinement).
+The 650M DPLM-2 exceeds the 150M across TM-score/RMSD under the same stochastic rollout settings, and MCTD-ME (de novo) is included for parity.
+The count is n=163n{=}163 (not $183$) because the evaluator filters by length (default $60\leq L\leq 512$), so out-of-range chains are skipped.
+As above, these de novo results should not be conflated with lead-optimization tables.
+Method TM-score to GT (mean $\pm$ std) [min–max] BB RMSD to GT (mean $\pm$ std) [min–max] Composite Reward (mean $\pm$ std) [min–max]
+DPLM-2_150Mstoch $0.656\pm 0.168$ [0.296–0.932] $11.015\pm 6.410$ [2.228–27.337] $0.333\pm 0.136$ [0.122–0.596]
+DPLM-2_650Mstoch $0.724\pm 0.160$ [0.299–0.937] $8.400\pm 5.819$ [1.852–30.628] $0.395\pm 0.137$ [0.123–0.618]
+DPLM-2_3Bstoch $0.345\pm 0.129$ [0.174–0.873] $40.777\pm 11.735$ [7.062–68.432] $142\pm 0.054 [0.0$73–0.371]
+MCTD-ME (MCTS_ME) $\mathbf{0.732\pm 0.140}$ [0.325–0.934] $\mathbf{8.331\pm 5.570}$ [2.209–29.407] $\mathbf{0.399\pm 0.126}$ [0.133–0.631]
+Table 7: One-shot forward folding (CAMEO subset, n=163n{=}163).
+Each model predicts a single structure per target from sequence (no refinement).
+We report backbone TM-score and backbone RMSD to ground truth, plus a composite reward (mean $\pm$ std across targets; min–max in brackets).
+Sampling configuration (for stochastic rows): --task folding --temperature 1.0 --sampling_strategy "annealing@2.2:1.0" --max_iter 150.
+MCTD-ME (MCTS_ME) uses the same rollout temperature/unmasking/sampling schedule, with planning via PH-UCT.
+Motif scaffolding (one-shot scaffolds).
+Table 8 reports per-motif one-shot results for DPLM-2 650M on the curated motif set of [42].
+Across 24 motifs (100 samples each), 17 motifs admit at least one feasible scaffold under our strict success criterion (377/2400 total successes).
+We report per-motif success rate together with motif RMSD, scTM, and pLDDT, establishing a transparent baseline and motivating our refinement/planning evaluations on the standard filtered set (motifs with $\geq$1 success).
+A.8.3 Evaluation metrics for motif-scaffolding baselines.
+scTM (structure-based TM-score).
+For each generated sequence we predict a backbone with ESMFold, then compute scTM by aligning this predicted structure to the corresponding reference PDB backbone using TM-align.
+We report the TM-score normalized by the reference chain length so that scores are comparable across proteins of different lengths.
+[…]
+Success criterion. A sample is counted as successful if it satisfies both scTM >0.8>0.8 and motif RMSD <1<1\,Å.
+Motif DPLM-2 650M RFDiffusion ProteinA FoldFlow
+[…]
+1BCF 1/100 8.424 0.807 69.5 0/100 5.714 0.316 77.1 0/100 9.427 0.260 0/100 7.471 0.194
+1PRW 65/100 0.990 0.837 83.3 0/100 6.281 0.743 90.2 0/100 9.905 0.181 0/100 7.607 0.164
+1QJG 0/100 5.029 0.688 78.9 0/100 5.216 0.406 81.9 0/100 5.033 0.142 0/100 4.700 0.164
+1YCR 10/100 0.909 0.679 67.1 0/100 1.695 0.151 78.6 94/100 0.342 0.914 1/100 2.744 0.364
+2KL8 54/100 1.007 0.838 72.9 0/100 5.374 0.645 77.4 0/100 5.453 0.877 0/100 10.127 0.106
+3IXT 33/100 0.683 0.720 72.8 0/100 4.776 0.085 68.1 100/100 0.159 0.970 0/100 7.965 0.128
+4JHW 0/100 9.540 0.672 61.7 0/100 5.770 0.171 70.3 0/100 12.654 0.126 0/100 12.927 0.138
+4ZYP 1/100 2.701 0.630 67.6 0/100 3.682 0.220 70.5 0/100 4.930 0.175 0/100 6.035 0.135
+5IUS 0/100 6.403 0.569 56.3 0/100 8.989 0.393 63.0 0/100 10.629 0.188 0/100 9.320 0.139
+5TPN 0/100 2.578 0.710 61.0 33/100 1.506 0.636 76.3 99/100 0.436 0.892 0/100 5.147 0.210
+5TRV_long 1/100 3.256 0.696 73.8 21/100 2.285 0.545 79.8 96/100 0.491 0.951 0/100 6.309 0.277
+5TRV_med 1/100 2.416 0.763 74.8 2/100 3.572 0.281 76.7 89/100 0.524 0.949 0/100 6.799 0.272
+5TRV_short 0/100 3.550 0.696 68.8 0/100 5.579 0.077 81.1 83/100 0.509 0.950 0/100 8.562 0.186
+5WN9 0/100 4.989 0.559 61.3 0/100 5.207 0.171 66.1 100/100 0.156 0.934 0/100 8.886 0.091
+5YUI 0/100 6.493 0.623 69.9 0/100 3.397 0.481 70.2 0/100 6.596 0.187 0/100 6.929 0.181
+6E6R_long 49/100 0.676 0.768 79.2 46/100 0.646 0.633 84.0 51/100 1.196 0.870 0/100 3.378 0.471
+6E6R_med 42/100 0.653 0.760 79.8 40/100 0.707 0.504 83.1 68/100 0.873 0.876 0/100 3.419 0.468
+6E6R_short 4/100 0.641 0.670 76.6 19/100 0.935 0.349 81.9 66/100 0.915 0.875 0/100 3.496 0.446
+6EXZ_long 35/100 1.117 0.772 74.8 56/100 0.938 0.639 84.7 100/100 0.180 0.924 0/100 3.807 0.347
+6EXZ_med 35/100 1.106 0.771 74.7 44/100 1.354 0.511 81.3 100/100 0.172 0.928 0/100 3.762 0.345
+6EXZ_short 3/100 2.136 0.663 65.3 37/100 1.476 0.503 82.9 100/100 0.177 0.926 0/100 4.182 0.329
+7MRX_128 13/100 2.484 0.735 69.8 19/100 2.204 0.476 72.9 100/100 0.165 0.943 0/100 5.044 0.241
+7MRX_60 12/100 1.530 0.772 70.8 21/100 3.026 0.357 73.0 100/100 0.174 0.941 0/100 5.305 0.208
+7MRX_85 18/100 1.986 0.809 71.4 17/100 2.666 0.380 72.2 100/100 0.174 0.942 0/100 5.291 0.201
+Table 8: One-shot motif scaffolding baselines (per-motif success and quality; Table 8).
+We evaluate each method as a one-shot scaffold generator on 24 curated motifs (following [42]), with 100 independent generations per motif.
+Succ.
+is the number of generations meeting our strict motif-preservation criterion.
+We report motif-aligned RMSD (lower is better), self-consistency TM-score (scTM; higher is better), and mean pLDDT confidence (higher is better).
+ProteinA and FoldFlow pLDDT values are unavailable in these runs and are reported as “–”.
+Metric DPLM-2 650M RFDiffusion ProteinA FoldFlow
+Motifs with $\geq 1$ success (/24/24) 17/24 12/24 16/24 1/24
+Total successes (/2400/2400) 377/2400 (15.7%) 355/2400 (14.8%) 1446/2400 (60.2%) 1/2400 (0.0%)
+Mean motif RMSD $\downarrow$ $2.971\pm 3.278$ $3.382\pm 2.583$ $2.969\pm 3.984$ $6.217\pm 2.610$
+Mean scTM $\uparrow$ $0.717\pm 0.129$ $0.403\pm 0.362$ $0.705\pm 0.340$ $0.242\pm 0.148$
+Mean pLDDT $\uparrow$ $70.9\pm 11.1$ $76.8\pm 11.9$
+Table 9: Motif scaffolding one-shot baselines (Table 8; summary over 24 motifs, 100 samples each).
+DPLM-2 650M numbers are from Table 8.
+RFDiffusion numbers follow the updated full-pipeline baseline (ProteinMPNN $\rightarrow$ ESMFold); mean RMSD is computed over motifs with defined RMSD.
+ProteinA and FoldFlow pLDDT values are unavailable in these runs and are reported as “–”.
+ProteinA is substantially stronger than FoldFlow on this motif-region evaluation: it achieves a $60.2\%$ overall success rate (1446/2400) and succeeds on 16/24 motifs, whereas FoldFlow only has 1/2400 successes (1/24 motifs with any success).
+Consistent with this, ProteinA attains much higher mean scTM ($0.705$ vs.
+$0.242$) and lower mean motif RMSD ($2.969$ vs.
+$6.217$), indicating markedly better motif fidelity and global structural agreement under our strict criterion.
+Compute note.
+The baseline numbers above are from 100 independent one-shot samples per motif.
+In contrast, in our main refinement/planning experiments we typically run only 25 search iterations per motif/target; despite this much smaller budget, the planner already surfaces many successful scaffolds in practice, demonstrating that multi-step guided search can identify high-quality candidates efficiently.
+[…]
+[…]
+$\rho\;\triangleq\;\tfrac{t}{T}\in[0,1],\qquad B(s,a)\;\triangleq\;w_{\mathrm{ent}}\,U_{\mathrm{ent}}(s,a)\;+\;w_{\mathrm{div}}\,U_{\mathrm{div}}(s,a),$
+where $t$ is the global selection step, $T$ is the total budget, $U_{\mathrm{ent}},U_{\mathrm{div}}\in[0,1]$, and the nonnegative weights are chosen so that $B(s,a)\in[0,1]$ for all $(s,a)$.
+We apply a power decay to the bonus:
+$B_{\mathrm{decay}}(s,a;\rho)\;=\;\bigl(B(s,a)\bigr)^{\,1-\rho}.$
+The selection score is then
+$\text{P}\mathcal{H}\text{-UCB}\text{-ME}^{\mathrm{decay}}(s,a)\;=\;Q(s,a)\;+\;c_{p}\,\frac{\sqrt{\log N(s)}}{1+N(s,a)}\;\pi_{\mathrm{cons},\tau}(a\mid s)\;\bigl(B(s,a)\bigr)^{\,1-\rho}.$
+[…]
+Notes.
+[…]
+A.10 Experimental configuration and hyperparameters.
+Search setup and hyperparameters.
 Component Symbol Setting(s) used
@@
 Hyperparameters are fixed for the main comparisons; only the presence/number of experts differs by mode.
-A.5 Runtime and Efficiency
+For equitable assessment, we compare all variants under matched search and decoding budgets.
+Unless otherwise specified, every method uses the same DPLM-2 decoding configuration (deterministic unmasking with argmax sampling) and identical reward evaluation.
+In each table/figure, competing methods share the same rollout budget and selection rule; only the availability of diffusion experts differs across modes (random_no_expert, single_expert, multi_expert).
+We conducted only light tuning around a small grid and then fixed a single setting for all reported comparisons.
+Models, experts, and baselines (full details for §5.1).
+We instantiate MCTD-ME with masked discrete diffusion as the rollout engine and vary the proposal experts by task.
+Sequence experts are pre-trained DPLM-2 models [36] of different capacities (150M, 650M, 3B).
+For motif scaffolding we consider structure-aware experts (RFDiffusion [39], ProteinA [9], FoldFlow [4]) both as single-expert baselines and as members of a multi-expert ensemble.
+All experts propose from the same masked input; a single composite critic scores candidates.
+Baselines are task-specific: Folding (seq $\to\!$ struct) uses one-shot ESMFold [18] on the native sequence (RMSD/TM vs.
+ground truth), while our planners propose structure-token edits via DPLM-2 conditioned on the sequence and compare directly to the reference backbone.
+Inverse folding (struct $\to\!$ seq) uses one-shot DPLM-2 (150M) conditioned on backbone tokens; MCTD-ME performs masked amino-acid fills given the backbone.
+Motif scaffolding (motif $\to\!$ scaffold) uses one-shot DPLM-2 (150M) co-generating sequence/structure with motif residues frozen; MCTD-ME edits scaffold positions (sequence and/or structure, depending on expert) while preserving the motif.
+In this study, we frame evaluation as lead optimization: starting from a one-shot initial lead and refine it.
+We emphasize that our goal is not to propose a new one-shot SOTA generator/backbone.
+Instead, MCTD-ME is a test-time sampling/guidance layer that improves a fixed initial lead via discrete masked edits under a fixed evaluation budget.
+Accordingly, we focus on Final and especially $\Delta$ (Final–Baseline) under equal compute; stronger backbones can be incorporated as additional experts within the same planning loop.
+We compare MCTD-0 (random fill; no experts), MCTD-1 (single proposal expert; separate runs per expert), and MCTD-ME (multi-expert ensemble).
+We also include two multi-expert ablations on the CAMEO subset (n=183n{=}183), keeping the initial leads, critics, and wall-clock budget fixed: (i) Sampling (no MCTS)—experts propose directly on the pLDDT-derived mask; we score with the composite reward and report the per-target best (no selection/backprop); and (ii) UCT only—MCTS with 2 UCB selection [13] using $\mathrm{UCB}=Q(s,a)+c_{p}\sqrt{\tfrac{\log N(s)}{N(s,a)}}$, omitting the uncertainty/novelty bonuses used in $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$.
+Sampling isolates the benefit of structure-aware masking and expert diversity but underperforms MCTD-ME on recovery-style objectives and shows higher variance; UCT recovers part of the planning benefit over Sampling yet remains below MCTD-ME due to reduced proposal diversity.
+Due to limited space, we defer additional baselines and experiments to Appendix A.8.
+A.11 Runtime and efficiency.
 The enhanced performance comes at the cost of additional computation.
@@
 Times include all steps (masking, diffusion rollouts, scoring, and tree updates).
-A.6 Generality
+A.12 Generality.
 Although our evaluation focused on the inverse folding problem, the MCTD-ME framework is inherently general and model-agnostic.